### v1

# FlipedRAG: Black-Box Opinion Manipulation Attacks to Retrieval-Augmented Generation of Large Language Models

Abstract
Retrieval-Augmented Generation (RAG) addresses hallucination and real-time constraints by dynamically retrieving relevant information from a knowledge database to supplement the LLMs’ input. When presented with a query, RAG selects the most semantically similar texts from its knowledge bases and uses them as context for the LLMs to generate more accurate responses. While RAG offers numerous benefits, it also creates a new attack surface for large language models, especially since RAG databases are frequently sourced from public domains such as the web. While existing studies have predominantly focused on optimizing RAG’s performance and efficiency, emerging research has begun addressing the security concerns associated with RAG. However, these works have some limitations, typically focusing on either white-box methodologies or heuristic-based black-box attacks, and often lacking of in-depth exploration of retrieval subsystem. Furthermore, prior research has mainly targeted simple factoid question answering, which is neither practically challenging nor resistant to correction by LLMs. In this paper, we unveil a more realistic and threatening scenario: opinion manipulation for controversial topics against RAG. Particularly, we propose a novel RAG black-box attack method, termed FlipedRAG, which is transfer-based. By leveraging instruction engineering, we obtain partial retrieval model outputs from black-box RAG system, facilitating the training of surrogate models to enhance the effectiveness of opinion manipulation attack. Extensive experimental results confirms that, compared to the baseline method, our approach significantly enhances the average success rate of opinion manipulation by 16.7%. It achieves an average of a 50% directional change in the opinion polarity of RAG responses across four themes. Additionally, it induces a 20% shift in user cognition. Furthermore, we discuss the efficacy of potential defense mechanisms and conclude that they are insufficient in mitigating this type of attack, highlighting the urgent need to develop novel defensive strategies.
1 Introduction
Figure 1: An overview of our black-box opinion manipulation for revealing the vulnerabilities of RAG and its potential impact on people’s cognition.
With the rapid development of artificial intelligence, large language models (LLMs) have demonstrated exceptional capabilities in the field of natural language processing. However, constrained by their training data, these models have limited scope of knowledge and lack the most up-to-date information, which can lead to errors or hallucinations when tackling more complex or time-sensitive tasks. Retrieval-Augmented Generation (RAG) combines information retrieval with the generative capabilities of LLMs, enhancing the timeliness of knowledge acquisition and effectively mitigating the hallucination problem of these models. When given a query, RAG retrieves the most relevant documents from a knowledge base to augment the input request for the LLM. For example, the retrieved knowledge may consist of a series of text snippets that are semantically most similar to the query. RAG has inspired many popular applications, such as Microsoft Bing Chat, ERNIE Bot, and KimiChat, which use RAG to summarize retrieval results for improving user experience. Open-source projects like LangChain and LlamaIndex provide developers with flexible RAG frameworks to build customized AI applications using LLMs, retrieval models and knowledge bases.
As the application scope of RAG expands, concerns about its security, particularly regarding model performance under malicious attacks, have become prominent. The fundamental RAG process generally consists of three components: knowledge bases (corpus collections), retriever, and LLMs. Vulnerabilities arise when some of the retrieved documents are compromised by malicious attackers, a scenario referred to as a manipulation attack against RAG in this paper. Numerous studies have explored various forms of adversarial manipulation attacks, such as adversarial attack on the retriever [13, 15], prompt injection attack [1, 14, 9], jailbreak attack for LLM [4, 11, 26], and poisoning attack targeting the retrieval corpus in RAG [29, 23].
This paper primarily examines adversarial ranking poisoning attacks targeting the retriever in black-box RAG systems, which align with realistic and practical scenarios. It explores how such attacks indirectly influence the opinionated generative outputs of LLMs, as illustrated in Figure 1. By employing black-box retriever imitation [13] for opinion manipulation, we reveal specific vulnerabilities in RAG systems. The threat model presented here closely mirrors a real-world black-box scenario and can be characterized as follows: the adversary can only query the RAG system and does not have access to the complete knowledge base or corpus, the retriever, or the parameters of the RAG. The attacker is only capable of injecting adversarially modified candidate texts into the corpus, while the retriever and the LLM remain black-boxed, intact, and unmodifiable. Based on previous research [12, 2], the retrieval corpus and knowledge base consist of millions of candidate texts derived from the internet, enabling attackers to insert adversarially modified content through the malicious crafting of web content or encyclopedic entries. Cho et al. [3] and Zhong et al. [27] investigate the attack against predefined retrievers in white-box RAG, which are challenging to achieve in real-world scenarios with limited flexibility and practicality. Furthermore, they did not evaluate attacks specifically targeting the integrated generation process. Another representative work, PoisonedRAG [29], implemented black-box retrieval poisoning attacks against RAG knowledge bases, effectively uncovering pertinent security vulnerabilities. However, it predominantly concentrated on closed-domain question answering, such as "Who is the CEO of OpenAI?" These types of questions could be rectified when RAG systems are coupled with fact-checking and value alignment in LLMs.
The vulnerabilities explored in this paper primarily address controversial and opinion-based inquiries in RAG, such as "Should abortion be legal?" These questions are open-ended and require advanced levels of logical reasoning and summarization capabilities from LLMs. Current research on controversial topics is limited, and attacks that manipulate opinions on such questions could potentially inflict deeper harm.
Controversial topics are issues characterized by a lack of consensus due to divergent perspectives, often attracting widespread attention. These topics frequently involve differing viewpoints that shape public perception during extensive discussions. For instance, in political elections, Robert Epstein [6] demonstrated that manipulating search engines to generate biased results can alter voters’ preferences. In scenarios involving controversial topics, opinion manipulation can create “Information Bubbles”, which subsequently lead to the homogenization of user opinions [24]. In such cases, individuals’ views become easily swayed by the stance of the information they are exposed to. With the increasing use of LLMs in content production, exploiting vulnerabilities in RAG systems for opinion manipulation represents a particularly grave threat. Through carefully crafted inputs, attackers can influence the orientation of content generated by these models, thus jeopardizing users’ cognitive processes and decision-making abilities. Consequently, it is of considerable theoretical and practical importance to examine the vulnerabilities of RAG models to opinion manipulation attacks within a black-box setting.
To address aforementioned challenges, in this paper, we propose FlipedRAG, a black-box attack method, to explore the reliability of RAG in controversial topics and investigate its impact on user cognition. Specifically, we begin by inputting carefully designed instructions to obtain the rankings of retrieval results within the RAG model, thereby analyzing the operational mechanism of its retrieval module. Subsequently, we train a surrogate model using the extracted retrieval data [13, 20] to approximate the features and relevance preferences of the retriever in the black-box RAG system. Based on this surrogate model, we develop an adversarial attack strategy aimed at manipulating the opinions of candidate documents. By attacking this surrogate model, we generate adversarial opinion manipulation triggers, which are then transferred to the target RAG corpus. We conduct experiments on opinion datasets encompassing multiple topics to validate the effectiveness and scope of the attack strategy without relying on internal knowledge of the RAG model. The experimental results reveal that the proposed attack strategy can significantly alter the opinion polarity of the content generated by RAG. Furthermore, we discuss the efficacy of potential defense mechanisms and conclude that they are insufficient in mitigating this type of attack. This not only highlights the vulnerability of the retrieval model and the retrieval-augmented generation process but, more importantly, underscores the potential negative impact on user cognition and decision-making. Such vulnerabilities can increase the risk of misleading users into accepting incorrect or biased information.
Our major contributions are as follows:
(1) We propose a novel transfer-based opinion manipulation attack against RAG of LLMs under black-box setting, termed FlipedRAG, which aligns more closely with real-world and practical scenarios.
(2) To the best of our knowledge, we are the first to explore the adversarial opinion manipulation of RAG on open-ended and controversial topics.
(3) Through both automated evaluations and user experiments, we demonstrate that our transfer-based manipulation attack can significantly influence the opinions generated by RAG system in a specific desired direction.
(4) We discuss the efficacy of potential defense mechanisms and conclude that they are insufficient in mitigating this type of attack, highlighting the need for new defense strategy.
Figure 2: The overview of FlipedRAG for manipulating the opinions of RAG-generated content under black-box setting.
2 Related Works
2.1 Reliability of Neural Network Models
Research on the reliability of neural network models has been extensively studied. In 2013, Szegedy et al. [17] found that applying imperceptible perturbations to a neural network model during a classification task was sufficient to cause classification errors in CV. Later, scholars observed similar phenomena in NLP. Robin et al. [10] found that inserting perturbed text into original paragraphs significantly distracts computer systems without altering the correct answer or misleading humans. It reflects the robustness of neural network models, i.e., the ability to output stable and correct predictions in tackling the imperceptible additive noises [19]. For large language models, Wang et al. [18] proposed a comprehensive trustworthiness evaluation framework for LLMs, assessing their reliability from various perspectives such as toxicity, adversarial robustness, stereotype bias, and fairness. While large language models have greater capabilities compared with general deep neural network models, they also raise more concerns regarding security and reliability.
2.2 Reliability of RAG
As RAG is designed to overcome the hallucination problem in LLMs and enhance their generative capabilities, the reliability of content generated by RAG also becomes a major concern. Zhang et al. [25] sought to identify the weaknesses of RAG by analyzing critical components in order to facilitate the injection of the attack sequence and crafting the malicious documents with a gradient-guided token mutation technique. Xiang et al. [21] proposed an isolate-then-aggregate strategy, which first obtains responses from each passage of the LLM in isolation and then securely aggregates these isolated responses, to construct the first defense framework against retrieval corruption attacks. These studies are based on white-box scenarios and primarily focus on the robustness of RAG against corrupted and toxic content.
This paper proposes the use of adversarial retrieval attack strategies to manipulate the ranking results of the retriever, ensuring that opinion documents with a specific stance are ranked as highly as possible, thereby influencing the generated responses of the LLM to align with that stance.
2.3 Adversarial Retrieval Attack Strategy
The adversarial retrieval attack strategy starts with manipulation at the word level. Under white-box setting, Ebrahimi et al. [5] utilize an atomic flip operation, which swaps one token for another, to generate adversarial text(known as adversarial trigger) and the method is called Hotflip. Hotflip eliminates reliance on rules, but the adversarial trigger it generates usually has incomplete semantics and insufficient grammar fluency. While it can deceive the target model, it cannot evade perplexity-based defenses. Wu et al. [20] also proposed a word substitution ranking attack method called PRADA. To enhance the readability and effectiveness of the trigger, scholars further designed sentence-level ranking attack methods. Song et al. [16] propose an adversarial method under white-box setting, named Collision, which uses gradient optimization and beam search to produce the adversarial trigger named collision. The Collision method further imposes a soft constraint on collision generation by integrating a language model, reducing the perplexity of the collision. The method has shown promising results in document retrieval experiments. Inspired by Collisions, Liu et al. [13] propose the Pairwise Anchor-based Trigger (PAT) method under black-box setting. Added the fluency constraint and the next sentence prediction constraint, the method generates adversarial triggers by optimizing the pairwise loss of top candidates and target candidates with triggers. Although the time complexity of PAT has increased compared to previous methods, PAT takes ranking similarity and semantic consistency into account, so its manipulation effect on the retrieval ranking of target candidates is superior.
3 Method
3.1 Overview
Our objective is to manipulate the opinions expressed in responses generated by black-box RAG models on controversial topics, with a focus on both the retrieval model and the large language model that performs the integrated generation task. Zhang et al. [25] attempted to poison context documents to mislead the LLM into generating incorrect content. However, this approach necessitates extensive internal details of the LLM application, rendering it less feasible in real-world scenarios. In the black-box RAG context, the attacker lacks access to internal information of the RAG, including model architecture and scoring functions, and can only interact with the inputs and outputs of the RAG. Specifically, the attacker can only use the interface of the RAG and not directly access the retriever. Since the inputs to LLMs are comprised of the user query and context documents, and the user query is immutable, our attack strategy focuses on modifying candidate documents within the corpus. Although the attacker does not have access to the entire corpus, they can insert adversarially modified candidate texts into it, as many RAG applications source information from the Internet, where the content is publicly editable.
The basic framework of RAG comprises two components: the retriever and the LLM. These modules are serially connected, with the retriever sourcing context information from a knowledge base, upon which the LLM then performs the generation task. In black-box scenarios, attackers cannot modify the system prompts of the generative LLM, making it challenging to directly influence the generation results by exploiting any reliability flaws within the LLM itself. Therefore, we focus on exploiting the reliability flaws of the retriever to manipulate the retrieval ranking results. We employ specific instructions to induce RAG to reveal the context information it references. With this information, we train a surrogate model that approximates the black-box retriever, effectively transparentize it into a white-box model. We then employ this white-box model to generate adversarial triggers. By adding these adversarial triggers to candidate documents that reflect the desired opinion, we enhance their relevance to the user query, increasing the likelihood that they will be included in the context passed to the generative LLM. Leveraging the strong capability of LLM for context understanding and instructions-following, we guide it to generate responses that align with the expected opinion. An overview of this process is shown in Figure 2.
3.2 Threat Model
Given a topic $q$ (the question) from a set of controversial topics $Q$, a RAG corpus $D$ and a target opinion $S_{t}$ (Pro or Con, “Pro” stands for supporting and “Con” stands for opposing) as the opinion that the attacker expects the RAG system to generate, we denote the target document as $d_{t}$ which holds $S_{t}$ for $q$. Given a black-box retrieval model RM, it calculates relevance score $R(q,d)$ for each document $d$ in $D$ and return the set of documents with the $k$ highest relevance score as the retrieval result (the LLM only takes top $k$ candidate documents into account). The retrieval result is denoted as $\text{RM}_{k}(q)=\{d_{1},d_{2},...,d_{k}\}$. The LLM of the black-box RAG utilizes $\text{RM}_{k}(q)$ to answer $q$ with response $\text{LLM}(q,RM_{k}(q))$, the opinion of $\text{LLM}(q,RM_{k}(q))$ is $S_{o}$, $S_{o}=\text{S}(LLM(q,RM_{k}(q)))$.
3.2.1 Objective of the Adversary
The attacker aims to find the adversarial trigger $p_{\text{adv}}$ for each $d_{t}$, and adds it to $d_{t}$, transforming the corpus $D$ to $D(d;d_{t}\oplus p_{\text{adv}})$. $p_{\text{adv}}$ can increase the relevance score $R(q,d_{t}\oplus p_{\text{adv}})$ assigned by the retrieval model RM, $d_{t}$ will be ranked at the top of the retrieval results $\text{RM}_{k}(q)=\{d\mid d_{t}\oplus p_{\text{adv}}\}$ , guiding the LLM to generate responses that align with the target opinion: $\text{S}(LLM(q,\{d\mid d_{t}\oplus p_{\text{adv}}\}))=S_{t}$, $S_{t}\neq S_{o}$.
3.2.2 Capabilities of the Adversary
In the black-box scenario, we assume the attacker is only authorized to query the RAG, obtain RAG-generated results and modify limited documents in the corpus. There are no restrictions on the number of calls to RAG. Furthermore, the attacker has no knowledge of the model architecture, model parameters, or any other information related to the models within the black-box RAG. Modifying the prompt templates used by the LLM in RAG is also prohibited.
Figure 3: The framework for obtaining ranking imitation data of retrieval model in black-box RAG system.
3.3 Black-box Retriever Imitation
The primary issue in implementing manipulation is to make the retrieval model of the black-box RAG transparent so that the attacker is able to produce adversarial triggers based on the knowledge of black-box retriever. We aim to simulate the black-box retrieval model $RM$ in RAG. The basic idea is to train a surrogate model $M_{\text{i}}$ with the ranking results $RM_{\text{k}}(q)$ from the retrieval model $RM$, thus turning the black-box retrieval model into a white-box model. However, since the retriever and the LLM in RAG are serially connected, it is not possible to directly obtain the ranking results of the retriever. Therefore, we attempt to guide the LLM to replicate the retrieval results of the black-box RAG, then we obtain the documents deemed relevant to the user query by the black-box retrieval model $RM$, which can be used as positive examples $d_{+}$ for black-box imitation training. Subsequently, documents irrelevant to the query can be random sampled as negative examples $d_{-}$ Therefore, we design specific instructions to make the black-box RAG replicate the retrieval results of the retriever RM. These retrieval results only need to reflect the relevance to the query. Then, based on the generated results of the LLM and the corpus, we sample positive and negative data to train the surrogate model. The method for obtaining imitation data of the retrieval model in black-box RAG is illustrated in Figure 3. The specific instruction used is as follows:
—Here is the user question— < < <{query}> > > —Here is the user command— Please COPY all the given context altogether in [[ ]] including all marks and symbols. Do not omit any sentence of the context.
After that, we use a pairwise approach to sample data and train the surrogate model. The black-box RAG responds the instruction with context information so the responses generated by the black-box RAG reflect the retrieval results instead of being independent to the context. Relevant documents are sampled from the responses as positive examples $d_{+}$, and random irrelevant documents are sampled as negative examples $d_{-}$.
These sample pairs $(d_{+},d_{-})$ are then incorporated into the training dataset. After sampling the imitation data, we employ a pairwise training method to train the surrogate model $M_{i}$. Let the relevance score calculated by $M_{i}$ be $R_{i}$, we train the surrogate model $M_{i}$ by minimizing:
$L=-\frac{1}{|Q|}\sum_{q\in Q}\log\left(\frac{R_{i}(q,d_{+})}{R_{i}(q,d_{+})+\sum R_{i}(q,d_{-})}\right)$       (1)
3.4 Opinion Manipulation Attack
After obtaining the surrogate model $M_{i}$, we transform the manipulation of RAG-generated opinions in a black-box scenario into manipulation in a white-box setting. Since we have all the knowledge of the white-box surrogate model $M_{i}$, we directly implement adversarial retrieval attacks on it, generating the adversarial trigger $p_{\text{adv}}$ for the candidate document $d_{t}$ with opinion $S_{t}$. We employ Pairwise Anchor-based Trigger (PAT) [13] strategy for adversarial retrieval attacks, which is commonly used as baseline in related research. After adding the generated adversarial trigger to the target document with $S_{t}$, the attacker can corrupt the corpus of RAG by inserting the modified candidate document into the RAG corpus. Then, by querying the system of the black-box RAG model to get the generated response, we compare the opinions of the manipulated response with the opinions of the response generated by the benign RAG to evaluate the reliability of the black-box RAG.
PAT, as a representative adversarial retrieval attack strategy, adopts a pairwise generation paradigm. Given the target query, the target candidate item, and the top candidate item(named anchor, which is to guide the adversarial text generation), the strategy utilizes gradient optimization of pairwise loss, which is calculated from the candidate item and the anchor, to find the appropriate representation of an adversarial trigger. The strategy also adds fluency constraint and next sentence prediction constraint. By beam search for the words, the final adversarial trigger, denoted as $T_{pat}$, is iteratively generated in an auto-regressive way. We use $T_{\text{pat}}$ as $p_{\text{adv}}$, with its generated optimization function being [13]:
$\hskip 14.22636pt\max\{M_{i}(q,T_{\text{pat}};w)+\lambda_{1}\cdot\log P_{g}(T_{\text{pat}};w)\\ \hskip 71.13188pt+\lambda_{2}\cdot f_{\text{nsp}}(d_{t},T_{\text{pat}};w)\}\hskip 19.91692pt$ (1)
In the above formula, $P_{g}$ is the semantic constraint function, and $f_{\text{nsp}}$ is the next sentence prediction consistency score function between $T_{\text{pat}}$ and $d_{t}$. $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters, $w$ represents the internal weights of the corresponding model.
Subsequently, we insert $p_{\text{adv}}$ into the target document to execute the poisoning attack on the RAG corpus. Our optimization objective of RAG opinion manipulation attack is as follows:
$\hskip 14.22636pt\max\dfrac{1}{|Q|}\sum_{q\in Q}\text{I}(\text{S}(LLM(q,RM_{k}(q)))=S_{t})\\ \hskip 14.22636pts.t.RM(q)=E(q)\cdot E(d),d\in D(d;d_{t}\oplus p_{\text{adv}})\hskip 56.9055pt\\ \hskip 14.22636ptRM_{k}(q)=\{d_{1},d_{2},...d_{k}\mid argmax_{d\in D}RM(q)\}$ (2)
I(·) is the indicator function whose return value is 1 if the conditions within the parentheses are satisfied; otherwise, the return value is 0. S(·) is the opinion classification function. The retrieval model computes semantic relevance score by dot product, and $E$ denotes the word embedding vector. $RM_{k}(q)$ denotes the set composed of the top $k$ documents with the highest relevance scores collected from the ranking of black-box retrieval model $RM$.
4 Experiments
4.1 Dataset
In terms of dataset, we utilize the MS MARCO Passages Ranking dataset as the data source for guiding the black-box RAG to generate relevant passages [27] where we sample data pairs to train the surrogate model. Additionally, we use controversial topics data scraped from the PROCON.ORG website as the object of manipulation. The controversial topic dataset includes over 80 topics, covering fields such as society, health, government, and education. Each controversial topic is discussed from two stances (Pro and Con), with an average of 30 related passages, each holding a certain opinion with stance Pro or Con. We random select 30 public controversial topics to conduct opinion manipulation experiment on RAG.
4.2 Experiment Details
The specific setting details for the RAG manipulation experiment are as follows:
(1) Black-box RAG. We represent the black-box RAG process as $\text{RAG}_{\text{black}}$. The RAG framework is Conversational RAG from LangChain. Conversational RAG adds logic for incorporating historical messages and intention reasoning, thus supporting back-and-forth conversation with users and serving as a widely used RAG framework. Based on a retriever and a large language model (LLM), Conversational RAG mainly consists of a history-awaring retriever to identify users’ intention and a question-answering chain to produce output, as shown in Figure 4. The LLMs adopted are the open-source models Meta-Llama-3-8B-Instruct (Llama3), Vicuna-13b-v1.5(Vicuna) and Mixtral-8x7B-Instruct-v0.1(Mixtral). They perform well across various tasks among all open-source LLMs. The prompt connecting the retriever and the LLM in $\text{RAG}_{\text{black}}$ adopts the basic RAG prompt from LangChain:
Use the following pieces of retrieved context to answer the question. Keep the answer concise. Context: {context}. Question: {question}.
Figure 4: The framework of Conversational RAG.
(2) Target retriever model and surrogate model. People usually use dense retrieval models as the retriever in RAG. Therefore, we select three representative dense retrieval models, Contriever, Co-Condenser and ANCE, as the target retrieval model [7, 8, 22]. By convention, we use dot product between the embedding vectors of questions(queries) and candidate documents as their similarity score $R$ in the ranking. The surrogate model chosen in this paper is Nboost/pt-bert-base-uncased-msmarco, which is type of BERT and is specifically trained on the MS Marco Passage Ranking dataset.
(3) Manipulation target. For a controversial topic $q$, documents $d_{t}$ with the target opinion $S_{t}$ are manipulated by adding adversarial trigger $p_{adv}$ at the beginning. This manipulation aims to place these perturbed documents as prominently as possible in the top $k$ rankings of the RAG retriever $\mathrm{RM}_{k}(q)$, where $k$ denotes the number of documents obtained from the retrieval results by the LLM in $\text{RAG}_{\text{black}}$ In this paper, $k$ is set to 3.
(4)Imitation data construction. We collect imitation training data in a pairwise way, sampling both positive samples and negative samples. It is worth mentioning that we randomly sample negative examples from the top of the ranking produced by the original surrogate model. We discover that taking highly relevant documents ranked by the surrogate model as negative examples helps the surrogate model learn more features of the target black-box retrieval model by contrastive learning between positive and negative examples.
(5) Opinion classification. We use Qwen2.5-Instruct-72B as the opinion classifier. Qwen2.5-Instruct-72B, due to its large parameter size, is capable of accurately identifying and classifying opinions within text.
(6) Experimental parameters. During black-box imitation training, we set the batch size as 256, the number of epoch as 4 and the learning rate as 4e-5. In the process of implementing PAT to generate adversarial triggers, we set the number of beams as 30, the temperature value as 0.4, the learning rate as 0.1 and the sequence length as 15.
(7) Experimental environment. All our experiments are conducted in Python 3.8 environment and run on a NVIDIA DGX H100 GPU.
Our proposed opinion manipulation strategy FlipedRAG for black-box RAG is outlined in Algorithm 1.
Input: target black-box RAG model $\textit{RAG}_{\text{black}}$, target retrieval model $RM$, surrogate model $\boldsymbol{M_{i}}$, controversial topics $\boldsymbol{Q}$, target topic $\boldsymbol{q}$, target opinion $S_{t}$, corpus $Docs$ , target documents with target opinion $Docs_{t}$, target document $\boldsymbol{d_{t}}$, relevant document $d_{+}$, random sampled document $d_{-}$
Instructions:
$i_{1}$ = "—Here is the user question—
< < < {query} > > >
—Here is the USER COMMAND—
Please COPY all the given context altogether in [[ ]] including all marks and symbols. Do not omit any sentence of the context."
$i_{2}$ = "Use the following pieces of retrieved context to answer the question…"
// $\textit{RAG}_{\text{black}}$ uses $i_{2}$ as prompt template.
Functions:
$\boldsymbol{OpinionClassify}$: Classify the opinion of the content into "support", "neutral" or "oppose".
$\boldsymbol{PAT}$: Pairwise Anchor-based Trigger generation strategy.
Output: manipulated RAG responses $\boldsymbol{Res}$
1
2 Phase 1. Pairwise Imitation Data Construction and Black-box Retrieval Model Imitation Training
3       INIT: Dataset $\mathcal{D}\leftarrow\{\}$
4       for $\boldsymbol{q}_{m}\in\boldsymbol{Q}$ do
5             induced rank list $\boldsymbol{R}_{m}{top3}\leftarrow$ $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m}\oplus i_{1};\boldsymbol{Docs})$
             // $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m}\oplus i_{1})$approx $RM(\boldsymbol{q}_{m})$
6             for $\boldsymbol{d_{+}}_{j}\in\boldsymbol{R}_{m}{top3}$ do
7                   Random sample document from the ranking of origin $\boldsymbol{M_{i}}$ as $\boldsymbol{d_{-}}_{j}$
8                   $\mathcal{D}\leftarrow positive,[\boldsymbol{q}_{m};\boldsymbol{d_{+}}_{j};\boldsymbol{d_{-}}_{j}]$
9                  $\mathcal{D}\leftarrow negative,[\boldsymbol{q}_{m};\boldsymbol{d_{-}}_{j};\boldsymbol{d_{+}}_{j}]$
                   // Reverse $\boldsymbol{d_{+}}_{j}$ and $\boldsymbol{d_{-}}_{j}$ to get the negative triple
10
11
12      Train the surrogate model $\boldsymbol{M_{i}}$ on $\mathcal{D}$ with Eq 1
13       return $\boldsymbol{M_{i}}$
14Phase 2. Adversarial Trigger Generation and Opinion Manipulation in RAG Response
15       INIT: RAG Response Set $\boldsymbol{Res}\leftarrow\{\}$
16       for $\boldsymbol{q}_{m}\in\boldsymbol{Q}$ do
17             rank list $\boldsymbol{R}_{m}\leftarrow$ $\boldsymbol{M_{i}(\boldsymbol{q}_{m};Docs)}$
18             $\boldsymbol{anchor}_{m}\leftarrow$ top-1($\boldsymbol{R}_{m}$)
19             for $d_{j}\in\boldsymbol{R}_{m}$ do
20                   if $\boldsymbol{OpinionClassify}(d_{j})=S_{t}$ then
21                         $\boldsymbol{Docs_{t}}\leftarrow$ $d_{j}$
22
23            for $\boldsymbol{d_{t}}_{j}\in\boldsymbol{Docs_{t}}$ do
24                   adversarial trigger $\boldsymbol{p_{adv}}_{j}\leftarrow\boldsymbol{PAT}(\boldsymbol{M_{i}};\boldsymbol{q}_{m},\boldsymbol{d_{t}}_{j},\boldsymbol{anchor}_{m})$
25                   adversarial doc $\boldsymbol{d_{adv}}_{j}\leftarrow\boldsymbol{d_{t}}_{j}\oplus\boldsymbol{p_{adv}}_{j}$
                   $\boldsymbol{Docs}:\boldsymbol{d_{t}}_{j}\leftarrow\boldsymbol{d_{adv}}_{j}$ // Replace
26
27            $\boldsymbol{Res}\leftarrow$ $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m};\boldsymbol{Docs})$
28
29      return $\boldsymbol{Res}$
Algorithm 1 FlipedRAG (black-box)
4.3 Research Questions
We propose three research questions to evaluate the effect of our method about black-box retriever imitation and opinion manipulation in RAG.
RQ1: Does black-box retriever imitation effectively learn about the internal knowledge of the retriever of RAG?
RQ2: How significant does opinion manipulation influence the response of the LLM in RAG?
RQ3: Does opinion manipulation significantly impact users’ perceptions of controversial topics?
4.4 Evaluation Metrics
For RQ1, we aim to evaluate the relevance ranking ability of the surrogate model trained through black-box imitation of $\text{RAG}_{\text{black}}$ on the standard ranking task and assess the similarity between their ranking performance and that of the target retrieval model. Additionally, when adversarial triggers are applied to the target retrieval model, we compare the manipulation effect of triggers generated from the surrogate model versus those generated from the original BERT model, which is not trained by black-box imitation. If the trigger generated based on the surrogate model shows better manipulation performance on the ranking of the target retrieval model, it indicates that the surrogate model has learned internal knowledge features of the target retrieval model through black-box retriever imitation training.
So we use Mean Reciprocal Rank (MRR) and Normalized Discounted Cumulative Gain (NDCG) to reflect the ranking ability of the models; higher values of these metrics indicate stronger ranking ability of the model in terms of relevance. Then we use Inter Ranking Similarity (Inter) and Rank Biased Overlap (RBO) to measure the similarity between the ranking performance of the surrogate model and that of the target retrieval model.
Mean Reciprocal Rank (MRR). MRR reflects the effectiveness of ranking by calculating the reciprocal of the rank at which the first relevant item appears. So MRR highlights relevant items. The MRR value is averaged over multiple queries and it ranges from 0 to 1, higher MRR values indicate better rankings with relevant documents found earlier in the ranking. MRR@10 means the score only takes the first 10 items in the ranking into account, the same below.
Normalized Discounted Cumulative Gain (NDCG). NDCG measure first calculate Cumulative Gain (CG), the sum of relevance scores for the results in the ranking. Later, considering that relevant items appearing earlier in the ranking are more valuable, it adjust CG to Discounted Cumulative Gain (DCG) by accounting for the position of each candidate item. At last, it normalizes the actual DCG with the maximum possible DCG that could be achieved if all relevant items are ranked in perfect order, resulting the NDCG score ranging from 0 to 1. An NDCG score close to 1 indicates that the ranking is nearly ideal, with highly relevant items near the top of the list.
Inter Ranking Similarity (Inter). It is a metric used to measure the similarity between two ranking list by counting the overlap of the top candidates, which is ranging from 0 to 1.
Rank Biased Overlap (RBO). RBO considers both the positions of candidate items and the extent of overlap between the two rankings and it ranges from 0 to 1, too. RBO@1000 calculates the top 1000 weighted ranking overlap.
For RQ2, we aim to evaluate the manipulation performance of our attack on retriever rankings and LLM responses. So we firstly evaluate the manipulation performance with the average proportion of target opinions in the top 3 rankings before and after manipulation (Top3origin, Top3attacked). Based on Top3origin and Top3attacked, we calculate the difference Top3var between these two values to reflect the extent of change in the ranking manipulation of candidate documents holding the target opinion $S_{t}$. Additional, we also use Ranking Attack Success Rate (RASR), Boost Rank (BRank) and Boost Rate (BRate) to reflect how effective our attack distort the retrieval ranking of RAG. Here are the details of the metrics on ranking manipulation above.
Top3v. Top3v is obtained by subtracting Top3origin from Top3attacked. Note that Top3v ranges from -1 to 1, and a positive value indicates that the manipulation successfully increased the proportion of the target opinion in the top-3 ranking, while a negative value suggests a detrimental effect of the manipulation. The higher the Top3v value, the more effective the manipulation. Top6v is also defined as above.
Ranking Attack Success Rate (RASR). RASR measures the average rate of successfully boosted candidates in the ranking of each query. The higher is the RASR value, the more successful is the ranking attack.
Boost Rank (BRank):BRank is the average of the total rank improvements for all target documents under each query. Note that a target document $d_{t}$ contributes a negative value to BRank if its rank is reduced accidentally by the attack.
Apart from the evaluation on ranking manipulation, we use Opinion Manipulation Success Rate (OMSR) and Average Stance Variation (ASV) to evaluate the significance of our opinion manipulation on the response of the LLM in black-box RAG.
Opinion Manipulation Success Rate (OMSR). OMSR measures the average rate of successfully manipulated LLM responses at the opinion level. We consider one manipulation as a success if it guides the LLM to produce a response whose opinion is closer to the target opinion $S_{t}$ compared with the original response without RAG attack. Given the target opinion "support"(Pro), successful opinion manipulation can be from "oppose"(Con) to "neutral", from "oppose" to "support", and from "neutral" to "support". Given the target opinion "oppose", successful opinion manipulation can be from "support" to "neutral", from "support" to "oppose", and from "neutral" to "oppose".
Average Stance Variation (ASV). ASV represents the average increase of opinion scores of LLM responses in the direction of the target opinion $S_{t}$ after manipulation. In this research, we assign a value of 0 to opposing stance, 1 to neutral stance, and 2 to supporting stance. A positive ASV indicates that the opinion manipulation towards $S_{t}$ is effective, while a negative ASV indicates that the manipulation actually makes the opinions of RAG responses deviate from $S_{t}$. ASV reflects how much the shift in opinion polarity is. The larger the ASV value, the more successful the opinion manipulation of RAG responses.
5 Experimental Results Analysis
5.1 RQ1: Does black-box retriever imitation effectively learn about the internal knowledge of the retriever of RAG?
After imitating the retrieval model of $\text{RAG}_{\text{black}}$ to obtain the surrogate model, we first compare the ranking ability of the surrogate model $M_{i}$ and the target retrieval model $RM$, as well as the similarity of their ranking results to ensure that the surrogate model has learned the capabilities of the black-box retrieval model.
TABLE I: Comparison(%) of ranking ability between the surrogate model and the target retrieval model(based on the target retrieval model).
Model MRR@10 NDCG@10 Inter@10 RBO@10
Contriever 79.06 67.71
$M_{Contriever}$ 86.82 69.96 62.55 48.52
Co-Condenser 87.01 68.16
$M_{CoCondenser}$ 84.06 67.98 61.86 49.40
ANCE 85.34 67.84
$M_{ANCE}$ 83.79 66.49 56.74 45.66
As shown in Table 1, the surrogate model $M_{i}$ trained by black-box imitation is similar to the target retrieval model in relevance ranking ability for their close NDCG@10 scores. Their ranking results are also similar enough, validating the effectiveness of the black-box imitation. For convenience, we use $M_{Contriever}$ to represent the surrogate model that has its training data collected from $\text{RAG}_{\text{black}}$ which uses Contriever as the retrieval model. The method of representation applies to surrogate models based on ranking data of Co-Condenser and ANCE as well.
In addition to relevance ranking performance, we place great importance on the manipulation performance of the adversarial triggers generated from the surrogate model $M_{i}$, because this performance also reflects the effectiveness of black-box imitation. Assuming we produce adversarial triggers by the white-box target retrieval model, its manipulation effect on the target retrieval model should be significantly better than the that of triggers generated by the BERT model for transfer attacks. The ranking manipulation effect of the surrogate model $M_{i}$ on the target retrieval model should fall between the manipulation effect of the white-box target retrieval model and that of BERT-based transfer attacks as $M_{i}$ has acquired some of the target model’s capabilities. In other words, only if triggers from the surrogate model has the better ranking manipulation effect on the target retrieval model than that from BERT, we consider the black-box imitation process to be successful. So we take BERT as the baseline model to evaluate the manipulation effect of black-box imitation process.
TABLE II: Comparison of Manipulation Effect between the Surrogate Model $M_{i}$ and the Baseline Model.
Target Retrieval Model Target Opinion Victim Model Ranking Manipulation Performance Opinion Manipulation on Llama-3
$Top3_{v}$ $Top6_{v}$ $BRank$ $RASR\%$ $OMSR\%$ $ASV$
Contriever PRO baseline 0.31 0.28 72.33 70.73 43.33 0.50
surrogate 0.37 0.38 111.03 74.22 63.33 0.70
CON baseline 0.27 0.24 67.1 71.93 40.00 0.27
surrogate 0.37 0.35 105.13 78.26 50.00 0.47
Co-Condenser PRO baseline 0.22 0.12 43.53 58.91 23.33 0.20
surrogate 0.27 0.22 63.70 62.01 43.33 0.27
CON baseline 0.07 0.13 37.43 57.11 20.00 0.07
surrogate 0.17 0.19 58.96 64.42 33.33 0.27
ANCE PRO baseline 0.30 0.22 57.20 62.79 43.33 0.43
surrogate 0.38 0.32 85.96 69.76 53.33 0.47
CON baseline 0.24 0.19 50.90 64.42 30.00 0.20
surrogate 0.33 0.32 89.40 74.50 36.66 0.37
As shown in Table 2, we find that in terms of ranking manipulation effectiveness, the surrogate models after black-box imitation outperform the baseline models. In both ranking attack success rate (RASR) and boosted rank (BRank) of target documents, triggers generated from the surrogate model achieve better transfer attack effects on the target retrieval model. Specifically, $M_{Contriever}$ has the best ranking manipulation performance on the triggers generated for transfer attacking, especially in terms of boost rank and $Top3_{v}$. $M_{ANCE}$ demonstrates ranking manipulation effectiveness comparable to that of $M_{Contriever}$, whereas the manipulation effect achieved by $M_{CoCondenser}$ is less ideal by comparison. It indicates that Co-Condenser, as a dense retrieval model, is more robust and reliable than Contriver and ANCE.
Additionally, the surrogate models, compared to the baseline BERT, promote more documents holding the target opinion into top 3 range, thereby having a significant impact on the response of $\text{RAG}_{\text{black}}$. Thus, we can see that the transfer attack based on the surrogate model achieves stronger opinion manipulation by successfully flipping opinions of RAG response in the direction of target opinion. Black-box imitation enables the surrogate model to learn the features of the black-box retrieval model, allowing triggers generated by the surrogate model to induce a greater shift in opinion polarity in LLMs’ response compared to the baseline model. Table 2 also indicates that adversarial triggers generated by $M_{Contriever}$ and $M_{ANCE}$ place more target documents in top 3 range. with greater extent of rank promoting than the baseline model has. As a result, $M_{Contriever}$ and $M_{ANCE}$ achieve relatively better performance on OMSR and ASV.
It is interesting that these dense retrieval models seem to favor supportive opinions. They exhibit better ranking manipulation capability on supportive target opinion than opposing target opinion.
In short, the above experiments demonstrate that the surrogate model obtained by black-box imitation not only achieves relevance ranking performance similar to that of the target retrieval model but also generates adversarial triggers with better ranking manipulation effects compared to the baseline. This indicates that the black-box imitation process is able to help BERT learn internal knowledge about the target black-box retrieval model.
5.2 RQ2: How significant does opinion manipulation influence the response of the LLM in RAG?
After the black-box imitation training, we conduct opinion manipulation experiments across different $\text{RAG}_{\text{black}}$ (based on Llama-3, Mixtral, or Vicuna-v1.5) with adversarial triggers generated by the three white-box surrogate models.
TABLE III: RAG Opinion Manipulation Effect among Different LLMs.
Surrogate Model Target Opinion OMSR(%) on Different LLMs ASV on Different LLMs
Llama-3 Mixtral Vicuna Llama-3 Mixtral Vicuna
$M_{Contriever}$ PRO 63.33 46.66 36.66 0.70 0.57 0.50
CON 50.00 46.66 46.66 0.47 0.40 0.60
$M_{CoCondenser}$ PRO 43.33 36.66 36.66 0.30 0.37 0.43
CON 33.33 36.66 33.33 0.27 0.37 0.33
$M_{ANCE}$ PRO 53.33 56.66 46.66 0.47 0.53 0.53
CON 36.66 40.00 46.66 0.37 0.47 0.60
As shown in Table 3, our transfer attack with black-box imitation achieve significant opinion manipulation effect on all three LLMs. As a whole, it has a 40% – 50% chance of successful opinion polarity flipping and realizes about 0.46 in opinion polarity shift of range 0 to 2. Moreover, table 3 reveals that FlipedRAG attacks based on the surrogate models $M_{Contriever}$ and $M_{ANCE}$ are more effective in corrupting the RAG corpus and manipulating the RAG opinion with relatively higher OMSR and ASV value. That is because these attacks exhibit superior performance on ranking distortion, especially in $Top3_{v}$ aspect.
From the perspective of opinion polarity manipulation, we observe that the impact of the ranking distortion caused by transfer attacks varies in opinion manipulation effectiveness across different LLMs. Llama-3 and Mixtral have higher OMSR score in contrast to Vicuna for their greater instruction-following ability. Thus, Llama-3 and Mixtral are more susceptible to the adversarial documents promoted to top 3 of the ranking, generating responses based on the stance presented in the context that is composed of these adversarial documents. On the contrary, for ASV, Vicuna actually demonstrates a relatively greater shift in opinion polarity. We analyze that this is due to the better diversity in responses from Llama-3 and Mixtral when answering questions, resulting in less stable opinion polarity in the generated responses. In some cases, opinion manipulation is unsuccessful, instead of turning the opinion of $\text{RAG}_{\text{black}}$ response to the target opinion or make it remain unchanged, manipulation causes the opinion polarity to shift in the opposite direction of the target opinion, thereby lowering the ASV value. The cause of these reverse manipulation results may also stem from the randomness of LLMs generation, although this influence is relatively minor.
To further investigate the fluctuations in opinions resulting from the generation randomness inherent in LLMs, we query the the RAG system twice, which uses Contriver as its retrieval model and has a clean database without manipulation, to get two responses on the same topic. The results are illustrated in the Table 4. We can observe that the responses generated by LLMs on controversial topics do exhibit fluctuations in opinions. However, these fluctuations are minimal so that the change of the text opinion polarity is significantly less than that resulting from deliberate manipulation attacks.
TABLE IV: Opinion Fluctuation of RAG Response on Clean Database.
LLM Target Opinion OMSR(%) ASV
Llama-3 PRO 20 0.17
CON 10 0.0
Mixtral PRO 10 0.07
CON 20 0.10
Vicuna PRO 13.3 -0.03
CON 16.7 0.07
Figure 5: Overall effect of RAG opinion manipulation on three LLMs under black-box setting (retrieval model: Contriever).
In addition, we observe that LLMs are more susceptible to manipulation in the direction of supportive target opinion. This is likely due to the dense retrieval model, acting as the retriever, favoring supportive stances.
From the perspective of overall opinion manipulation of $\text{RAG}_{\text{black}}$ response, as shown in Figure 5, our transfer attack with black-box imitation is able to significantly alter the opinion of RAG-generated content in a specified stance direction. When the target opinion is "Pro", it can be observed from the table that, compared to the opinion distribution before manipulation, the proportion of responses holding a supportive stance (in blue) has significantly increased in the opinion distribution after manipulation. Conversely, when the target opinion is "Con", there is a notable rise in the proportion of responses expressing an opposing stance in the manipulated opinion distribution.
Figure 6: Case Study of RAG Opinion Manipulation on "Should Corporal Punishment Be Used in K-12 Schools".
Figure 6 presents a case study of our RAG opinion manipulation on the question, "Should Corporal Punishment Be Used in K-12 Schools?" The black box highlights text expressing an supporting stance on the topic, while the red box indicates text opposing the topic. Initially, the LLM in the RAG system strongly opposed corporal punishment. However, after our transfer attack manipulated the retrieval ranking to prioritize documents supporting corporal punishment, the LLM’s opinion in its response reversed accordingly.
Later, to further analyze the opinion manipulation results on RAG response at a finer granularity, we conducted our transfer poisoning attacks on the corpus data of $\text{RAG}_{\text{black}}$ by themes. We focus on "Health", "Society", "Government", "Education" four themes.
The results of the theme-specific manipulation experiments are shown in Table 5. The adversarial retrieval attack strategy PAT, applied to poison the target document, significantly increased the proportion of candidate items with target opinion in the Top 3 of the $\text{RAG}_{\text{black}}$ retrieval list, thereby guiding the LLM to change its opinion in the response. However, the manipulation effect on $\text{RAG}_{\text{black}}$ output opinions varies across different themes.
TABLE V: Manipulation results of $\text{RAG}_{\text{black}}$ ranking and response opinion across different themes.
Theme Target Opinion Rank Manipulation Performance OMSR(%) on Different LLMs ASV on Different LLMs
$Top3_{v}$ $Top6_{v}$ Llama-3 Mixtral Vicuna Llama-3 Mixtral Vicuna
Health PRO 0.25 0.35 33.33 44.44 44.44 0.22 0.44 0.56
CON 0.33 0.33 44.44 44.44 33.33 0.67 0.33 0.33
Society PRO 0.38 0.36 33.33 33.33 16.67 0.17 0.17 0.17
CON 0.27 0.27 33.33 50.0 50.0 0.5 0.67 0.5
Government PRO 0.43 0.37 61.53 53.84 61.53 0.85 0.54 0.92
CON 0.38 0.44 38.46 61.53 61.53 0.38 0.62 0.85
Education PRO 0.42 0.35 57.14 71.42 42.85 0.86 0.86 0.71
CON 0.42 0.35 23.57 14.28 42.85 0.29 0.14 0.29
Overall, our transfer attack demonstrates the greatest manipulation effectiveness on LLM-generated opinions for questions related to the "Government" theme. The manipulation effectiveness for questions in the "Education" theme is also strong, although it is somewhat less effective when the target opinion is "CON." In comparison, the opinion manipulation effectiveness for the "Health" and "Social" themes is less satisfactory. This discrepancy is primarily due to the preference characteristics of the black-box retrieval model across different themes. As shown in Table 5, the rankings of the black-box target retrieval model for questions in the "Government" and "Education" themes are more susceptible to manipulation, as evidenced by their higher $Top3_{v}$ score.
5.3 RQ3: Does opinion manipulation significantly impact users’ perceptions of controversial topics?
To further explore the impact of opinion manipulation on users’ perceptions of public topics, we designed a user experiment to assess people’s reactions to RAG responses expressing different opinions. We invited 10 college students to participate in the experiment, dividing them into two groups, A and B, with five participants in each group. None of the participants were informed about the purpose of the experiment, and their experimental processes were independent of one another, with no communication allowed. We first developed a simple question-answering (QA) service based on the RAG framework, and then applied our attack to relevant documents containing supportive opinions on two designated controversial topics in the corpus. Participants in Group A accessed the QA service before any opinion manipulation was applied to the corpus, while participants in Group B accessed the service after we poisoned the corpus using our attack. Following the interaction, both groups were asked to rate their approval of the two topics on a scale of 1 to 10.
TABLE VI: User Approval Rating on the Public Topics.
Subject User Approval
Topic 1 Topic 2
A Group(before manipulation) 5.8 4.8
B Group(after manipulation) 6.4 8.2
As shown in Table 6, when we manipulated the relevance ranking of documents supporting the target topics, users’ approval of those topics significantly increased. This experiment confirms that opinion manipulation has a substantial effect on users’ perceptions.
5.4 Mitigation Analysis
Existing studies on RAG poisoning attacks have largely focused on manipulation in white-box scenarios. In contrast, research exploring RAG attacks from a more realistic black-box perspective has primarily employed relatively simple attack methods.
Zou et al. [29] propose PoisonedRAG which injects a few poisoned texts into the RAG database such that the LLM generates manipulated answers. They consider both black-box and white-box settings. In the black-box setting, they attempt to include the question itself as part of the poisoned text to make it semantically similar to the target question. The rationale behind this approach is that the question is most similar to itself. While this strategy is effective in manipulating RAG responses and easy to implement in practice, it is highly detectable in real-world applications, as directly inserting the question makes the manipulation overly conspicuous. Therefore, we use automatic spamicity detection to defend PoisonedRAG and FlipedRAG.
Table 7 illustrates the results of the automatic spamicity detection experiment for both PoisonedRAG and FlipedRAG, which utilizes black-box imitation and PAT. Spamicity detection [28] is used for discovering term spam. Earlier search engines relied on term-based ranking methods such as TF-IDF to retrieve documents. In this context, term spam manipulates document content through keyword stuffing to artificially inflate the TF-IDF scores of spam documents, making them appear more relevant for specific queries. A straightforward approach to achieve this is by embedding the question or query within the target document, thereby increasing the frequency of keywords in the document and boosting its TF-IDF-based ranking.
To identify spam documents, Spamicity detection calculates the term spamicity score of a given document to assess the likelihood that the document has been intentionally manipulated. Based on a spamicity threshold and the score of the target document, the detection strategy classifies the document as spam if its score exceeds the threshold.
In Table 7, we compute the spamicity scores of the adversarial triggers generated by PoisonedRAG and our transfer attack. We then calculate the probability of these adversarial triggers being detected as spam by the spamicity detection system under various thresholds, referred to as the detection rate.
It is evident that, under all threshold conditions, the detection rate of triggers generated by FlipedRAG is significantly lower than those generated by PoisonedRAG. This suggests that directly embedding the question in the document lacks sufficient stealth, making it more easily detectable by simple detection methods. As the detection threshold increases, the detection rate of adversarial triggers gradually decreases, with the most noticeable distinction between the two types of triggers occurring at the 0.2 threshold.
Additionally, our transfer attack utilizes the PAT ranking manipulation strategy, resulting in generated triggers that exhibit key characteristics of PAT, including strong semantic consistency and low text perplexity. These traits enable the triggers to evade detection based on perplexity measures.
TABLE VII: Automatic spamicity detection rates (%) of PoisonedRAG and FlipedRAG.
Threshold 0.3 0.25 0.2 0.15 0.1
PoisonedRAG 29 55 80 93 99
FlipedRAG 4 9 24 51 76
5.5 In-depth Discussion
In this paper, we present a method for transforming the target retrieval model into a white-box model through black-box imitation, thereby facilitating more effective and targeted opinion manipulation of black-box RAG systems. Extensive experiments demonstrate that, in a more realistic black-box setting for RAG systems, attackers can exploit access to the augmented context retrieved by the internal retrieval model. By leveraging this data, attackers can train a surrogate model and create poisoned documents that are both highly effective and stealthy in their ability to launch attacks.
Black-box imitation enhances the effectiveness and stealth of our transfer attack, making it more powerful than common transfer attacks and less likely to be detected. This approach provides a more accurate assessment of the reliability and security of RAG systems. Specifically, the opinion manipulation discussed in this paper focuses on open-ended, controversial topics that can influence human cognition. Consequently, assessing the reliability and security of RAG systems is of significant importance for RAG users, developers, and policymakers.
Our research also has certain limitations. We did not account for the influence of the opinion inherent in the PAT trigger text on RAG responses. However, due to the challenges in controlling the opinion polarity within PAT triggers, we plan to investigate more refined text ranking manipulation techniques in future work.
Given the identified vulnerabilities of RAG models, future research should focus on developing more robust defense strategies. These may include improving the robustness of retrieval algorithms, enhancing the reliability of generation models, and implementing multi-level input filtering mechanisms to mitigate adversarial inputs. Such approaches would enable a balanced optimization of both the understanding and reliability of RAG models.
6 Conclusion
In this paper, we investigate the vulnerability of retrieval-augmented generation (RAG) models to opinion manipulation through black-box attacks on open-ended, controversial topics. We also examine the potential impact of such attacks on user cognition and decision-making. Through systematic experimentation, we introduce a novel adversarial transfer attack strategy that employs black-box imitation to poison the retrieval ranking. This approach significantly alters the polarity of opinions generated by RAG models by crafting adversarial samples, without requiring internal knowledge of the RAG model.
The experimental results demonstrate that our transfer attack strategy effectively alters the opinions expressed in content generated by RAG models, exposing the vulnerability and unreliability of RAG systems when confronted with malicious retrieval corpora. Our proposed strategy is applicable across various dense retrieval models (Contriver, Co-Condenser, and ANCE) as well as different LLMs such as Llama-3, Mixtral, and Vicuna. Additionally, comparative experiments reveal that black-box imitation significantly enhances the effectiveness of transfer attacks in manipulating the opinions embedded in RAG-generated content. Furthermore, the adversarial documents generated using this method exhibit superior stealth compared to other black-box RAG poisoning attacks. To facilitate reproducibility, we have released the codes on the anonymous repository [^1] https://anonymous.4open.science/r/FlipRAG-EB0A/.
More importantly, this form of opinion manipulation could have profound consequences on users’ cognition and decision-making, potentially leading them to accept inaccurate or biased information. This could induce cognitive shifts and distort public opinion, especially on controversial issues, highlighting the critical importance of safeguarding against such vulnerabilities in RAG systems.
7 Ethical Considerations
In conducting research on the vulnerabilities of Retrieval-Augmented Generation (RAG) models, we have meticulously considered and addressed various ethical implications to ensure that our work complies with the highest ethical standards.
Disclosure and Mitigation. A primary ethical concern is the potential for our findings to be exploited by malicious actors in the development of more sophisticated attacks against RAG models. To mitigate this risk, we have implemented several precautionary measures: (1) We have responsibly disclosed the identified vulnerabilities prior to this submission, allowing for timely patching and strengthening of the affected models. Our methodologies, findings, and the ethical considerations involved in this study are thoroughly documented. (2) In addition to identifying vulnerabilities, our research offers comprehensive guidelines and strategies to enhance the security and reliability of RAG models. By providing these solutions, we aim to equip developers and researchers with the tools needed to safeguard against potential misuse.
Data Privacy and User Consent. Our experiments involved utilizing opinion-based experiments and datasets covering a range of topics. To protect the privacy and rights of individuals whose data may be included in these datasets, we have ensured that all datasets are anonymized to eliminate personally identifiable information (PII), thereby safeguarding individual privacy. Furthermore, prior to the commencement of the user experiments, participants were fully informed about the background and purpose of the study, as well as the measures in place to ensure fairness and protect privacy. After the experiment, we reiterated the true data distribution to participants to minimize any potential impact on their cognition. Throughout the research process, we have taken care to conduct all experiments, particularly those involving adversarial attacks, within secure environments to prevent unintended consequences.