### v3

# FlippedRAG: Black-Box Opinion Manipulation Adversarial Attacks to Retrieval-Augmented Generation Models

Abstract.
Retrieval-Augmented Generation (RAG) enriches LLMs by dynamically retrieving external knowledge, reducing hallucinations and satisfying real-time information needs. While existing research mainly targets RAG’s performance and efficiency, emerging studies highlight critical security concerns. Yet, current adversarial approaches remain limited, mostly addressing white-box scenarios or heuristic black-box attacks without fully investigating vulnerabilities in the retrieval phase. Additionally, prior works mainly focus on factoid Q&A tasks, their attacks lack complexity and can be easily corrected by advanced LLMs. In this paper, we investigate a more realistic and critical threat scenario: adversarial attacks intended for opinion manipulation against black-box RAG models, particularly on controversial topics. Specifically, we propose FlippedRAG, a transfer-based adversarial attack against black-box RAG systems. We first demonstrate that the underlying retriever of a black-box RAG system can be reverse-engineered and approximated by enumerating critical queries, candidates, and answers, enabling us to train a surrogate retriever. Leveraging the surrogate retriever, we further craft target poisoning triggers, altering vary few documents to effectively manipulate both retrieval and subsequent generation, transferring the attack to the original black-box RAG model. Extensive empirical results show that FlippedRAG substantially outperforms baseline methods, improving the average attack success rate by 16.7%. Across four diverse domains, FlippedRAG achieves on average a 50% directional shift in the opinion polarity of RAG-generated responses, ultimately causing a notable 20% shift in user cognition. Furthermore, we actively evaluate the performance of several potential defensive measures, concluding that existing mitigation strategies remain insufficient against such sophisticated manipulation attacks. These results highlight an urgent need for developing innovative defensive solutions to ensure the security and trustworthiness of RAG systems.
RAG, Black-box Adversarial Attack, Opinion Manipulation
*Corresponding author
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Software and application securityccs: Information systems Information retrievalccs: Computing methodologies Natural language processing
1. Introduction
Retrieval-Augmented Generation (RAG) combines information retrieval with the generative capabilities of LLMs, enhancing the timeliness of knowledge acquisition and effectively mitigating the hallucination problem of these models (Gao et al., 2023; Zhao et al., 2024b).
As the application scope of RAG expands, concerns about its security are also increasing. The basic RAG process typically consists of three components: a knowledge base (corpus collection), a retriever, and an LLM. Vulnerabilities arise when some of the retrieved documents are compromised, or prompt are tampered by malicious attackers, a scenario referred to as a manipulation attack against RAG. Previous studies have explored various forms of adversarial manipulation attacks, such as adversarial attack on the retriever (Liu et al., 2023b, c), prompt injection attack (Cai et al., 2022; Liu et al., 2023a; Jain et al., 2023), jailbreak attack for LLM (Deng et al., 2023; Li et al., 2023; Zhao et al., 2024a), and poisoning attack targeting the retrieval corpus in RAG (Cho et al., 2024; Zou et al., 2024; Xue et al., 2024).
Figure 1. (a) ChatGPT can be induced through user prompts to fully replicate the context segments referenced in its responses. Based on this observation, we can transparentize the retriever of the black-box RAG model by enumerating critical queries and candidates. (b) An overview of our black-box opinion manipulation adversarial attacks to RAG models and its potential impact on user’s cognition.
However, previous studies (Zou et al., 2024; Xue et al., 2024; Cho et al., 2024) have primarily focused on attacks against RAG systems, which lack feasibility and have fundamental limitations. They mostly address white-box scenario or employ heuristic-based black-box attacks, without thoroughly analyzing the vulnerabilities within the retrieval stage. Furthermore, earlier research has mainly concentrated on factoid question answering tasks, such as questions like “Who is the CEO of OpenAI?” Those factual queries are less impactful, as RAG systems combined with fact-checking and value alignment mechanisms in LLMs are capable of rectifying these errors (Xiang et al., 2024; Tu et al., 2025).
Differently, the scenario explored in this paper primarily addresses controversial and opinion-based inquiries in RAG, such as “Should abortion be legal?” These questions are open-ended and require advanced levels of logical reasoning and summarization capabilities from LLMs. Current research on controversial topics is limited. In scenarios involving controversial topics, opinion manipulation can create “Information Bubbles”, which subsequently lead to the homogenization of user opinions (Yue et al., 2023). In such cases, individuals’ views become easily swayed by the stance of the information they are exposed to. Attacks that manipulate opinions on such questions could potentially inflict deeper harm (Epstein and Robertson, 2015). Through carefully crafted inputs, attackers can influence the orientation of content generated by RAG models, thus jeopardizing users’ cognitive processes and decision-making abilities.
Thus, this paper primarily investigates adversarial opinion manipulation targeting the retriever in black-box RAG systems, which align with realistic and practical scenarios. The threat model presented here can be characterized as follows: the adversary can only query the RAG system and does not have access to the complete knowledge base or corpus, the retriever, or the parameters of the RAG. The attacker is only capable of injecting limited adversarially modified candidate texts into the corpus, while the retriever and the LLM remain black-boxed, intact, and unmodifiable.
To address the aforementioned challenges, in this paper, we propose FlippedRAG, a black-box attack method, to explore the reliability of RAG in controversial topics and investigate its impact on user cognition, as shown in Figure 1. By employing black-box retriever imitation (Liu et al., 2023b) for opinion manipulation, we reveal specific vulnerabilities in RAG systems. Our motivation is based on the observation in Figure 1(a). When we queried the RAG on a specific topic and included in the user prompt an instruction requesting LLMs to output the original content of its referenced retrieval results, it would faithfully reproduce the exact text segments it referenced. The reproduced outputs precisely matched the corresponding content on the web pages linked in the reference citations and reflect the functionality of relevance matching. It allows us to design a retriever imitation method for black-box RAG to address the limitations of white-box attacks or heuristic-based attacks against such systems. Specifically, we utilize carefully designed instructions to obtain the retrieval results within the RAG model. Subsequently, based on the constructed pairs (query, candidate) from the retrieval list, we train a surrogate model using the extracted retrieval data (Liu et al., 2023b; Wu et al., 2022b) to approximate and transparentize the relevance preferences of the retriever in the black-box RAG system.
Based on this surrogate model, we develop an attack strategy aimed at manipulating the opinions of candidate documents. By attacking this surrogate model, we generate adversarial opinion manipulation triggers, which are then transferred to the target RAG corpus, as shown in Figure 1 (b). We conduct experiments on opinion datasets encompassing multiple topics to validate the effectiveness and scope of the attack strategy without relying on internal knowledge of the RAG model. The experimental results reveal that the proposed attack strategy can significantly alter the opinion polarity of the content generated by RAG. Furthermore, we discuss the efficacy of potential defense mechanisms and conclude that they are insufficient in mitigating this type of attack. This not only highlights the vulnerability of the retrieval model and RAG model but, more importantly, underscores the potential negative impact on user cognition and decision-making. Such vulnerabilities can increase the risk of misleading users into accepting incorrect or biased information.
Our major contributions are as follows:
(1) We propose FlippedRAG, a novel transfer-based opinion manipulation attack against black-box RAG systems. It realistically reflects practical threat scenarios, and to our knowledge, we are among the first to study adversarial opinion manipulation targeting open-ended and controversial topics in RAG-based models.
(2) We propose a new approach to extract context documents and constructing pseudo-relevant training data for surrogate retriever models in black-box settings. It highlights risks regarding context data leakage in LLM applications. Extensive automated evaluations combined with user studies confirm that our method effectively manipulates the opinions generated by RAG models, ultimately influencing user perception toward specifically targeted viewpoints.
(3) We thoroughly evaluate existing defense measures and find that they fail to effectively prevent or reduce the impact of FlippedRAG, underscoring the urgent need for novel and more robust mitigation strategies.
2. Related Works
2.1. Reliability of Neural Network Models
Research on the reliability of neural network models has been extensively studied. In 2013, Szegedy et al. (Szegedy et al., 2013) found that applying imperceptible perturbations to a neural network model during a classification task was sufficient to cause classification errors in CV. Later, scholars observed similar phenomena in NLP. Robin et al. (Jia and Liang, 2017) found that inserting perturbed text into original paragraphs significantly distracts computer systems without altering the correct answer or misleading humans. It reflects the robustness of neural network models, i.e., the ability to output stable and correct predictions in tackling the imperceptible additive noises (Wang et al., 2019). For LLMs, Wang et al. (Wang et al., 2023) proposed a comprehensive trustworthiness evaluation framework, assessing their reliability from various perspectives such as toxicity, adversarial robustness, stereotype bias, and fairness. While LLMs have greater capabilities compared with general deep neural network models, they also raise more concerns regarding security and reliability.
2.2. Adversarial Retrieval Attack Strategy
The adversarial retrieval attack strategy starts with manipulation at the word level. Under white-box setting, Ebrahimi et al. (Ebrahimi et al., 2017) utilize an atomic flip operation, which swaps one token for another, to generate adversarial text, named Hotflip. Hotflip eliminates reliance on rules, but the adversarial trigger it generates usually has incomplete semantics and insufficient grammar fluency. While it can deceive the target model, it cannot evade perplexity-based defenses. Wu et al. (Wu et al., 2022b) also proposed a word substitution ranking attack method called PRADA. To enhance the readability and effectiveness of the trigger, Song et al. (Song et al., 2020) propose an adversarial method under white-box setting, named Collision, which uses gradient optimization and beam search to produce the adversarial trigger. Collision method further imposes a soft constraint on collision generation by integrating a language model, reducing the perplexity of the collision. Futhermore, Liu et al. (Liu et al., 2023b) propose the Pairwise Anchor-based Trigger (PAT) generation under black-box setting. By adding the fluency and semantic consistency constraints, it generates adversarial triggers by optimizing the pairwise loss of top candidates and target candidates with triggers.
2.3. Reliability of RAG
Some studies have investigated the adversarial poisoning attacks on RAG models, including PoisonedRAG(Zou et al., 2024), BadRAG(Xue et al., 2024), Phantom(Chaudhari et al., 2024), and GARAG(Cho et al., 2024). The distinctive characteristics of these attack methodologies are summarized in the Table 1. “LLM Capability Utilization” refers to the motivation of leveraging LLMs’ capabilities to induce malicious outputs. “Depth” denotes the cognitive penetration depth of RAG attacks. Although both BadRAG(Xue et al., 2024) and Phantom(Chaudhari et al., 2024) target user perception, their manipulations remain confined to sentiment level without achieving deeper ideological persuasion.
As RAG is designed to overcome the hallucination problem in LLMs and enhance their generative capabilities, the reliability of content generated by RAG also becomes a major concern. Zhang et al. (Zhang et al., 2024) sought to identify the weaknesses of RAG by analyzing critical components in order to facilitate the injection of the attack sequence and crafting the malicious documents with a gradient-guided token mutation technique. Xiang et al. (Xiang et al., 2024) proposed an isolate-then-aggregate strategy, named RobustRAG, which first obtains responses from each passage of the LLM in isolation and then securely aggregates these isolated responses, to construct the first defense framework against retrieval corruption attacks. These studies are based on white-box scenarios and primarily focus on the robustness of RAG against corrupted and toxic content.
Table 1. Comparison of existing RAG attacks. DoS denotes denial-of-service.
Methods Query type Scenario LLM Capability Utilization Depth
PoisonedRAG (Zou et al., 2024) Closed-ended, Factoid White-box/Black-box Logical reasoning
BadRAG (Xue et al., 2024) Open-ended, Factoid, DoS White-box Logical reasoning Sentiment
Phantom (Chaudhari et al., 2024) Open-ended, Factoid, DoS White-box Instruction-following Sentiment
GARAG (Cho et al., 2024) Closed-ended, Factoid Gray-box Noise resistence
FlippedRAG Controversial, Open-ended Black-box Induction, logical reasoning Perspectives
This paper proposes the use of adversarial retrieval attack to manipulate the retrieval process in black-box RAG, ensuring that corrupted documents with a specific opinion are ranked as highly as possible, thereby inducing the LLM responses to align with that opinion on controversial topics.
3. Methodology
3.1. Overview
Our objective is to manipulate the opinions expressed in responses generated by black-box RAG models on controversial topics, with a focus on both the retrieval model and the LLM that performs the integrated generation task. Zhang et al. (Zhang et al., 2024) attempted to poison context documents to mislead the LLM into generating incorrect content. However, this approach necessitates extensive internal details of the LLM application, rendering it less feasible in real-world scenarios. In the black-box RAG context, the attacker lacks access to internal information of the RAG, including model architecture and scoring functions, and can only interact with the inputs and outputs of the RAG. Specifically, the attacker can only use the interface of the RAG and not directly access the retriever. Since the inputs to LLMs are comprised of the user query and context documents, and the user query is immutable, our attack strategy focuses on modifying candidate documents within the corpus. Although the attacker does not have access to the entire corpus, they can insert adversarially modified candidate texts into it, as many RAG applications source information from the Internet, where the content is publicly editable.
The basic framework of RAG comprises two components: the retriever and the LLM. These modules are serially connected, with the retriever sourcing context information from a knowledge base, upon which the LLM then performs the generation task. In black-box scenarios, attackers cannot modify the system prompts of the generative LLM, making it challenging to directly influence the generation results by exploiting any reliability flaws within the LLM itself. Therefore, we focus on exploiting the reliability flaws of the retriever to manipulate the retrieval ranking results. We employ specific instructions to induce RAG to reveal the context information it references. With this information, we train a surrogate model that approximates the black-box retriever, effectively transparentize it into a white-box model. We then employ this white-box model to generate adversarial triggers. By adding these adversarial triggers to candidate documents that reflect the target opinion, we enhance their relevance to the user query, increasing the likelihood that they will be included in the context passed to the generative LLM. Leveraging the strong capability of LLM for context understanding, summarization and instructions-following, we guide it to generate responses that align with the expected opinion. An overview of this process is shown in Figure 2. The proposed method is outlined in Algorithm 1 in Appendix A.1.
Figure 2. The overview of FlippedRAG for manipulating the opinions of RAG-generated content under black-box setting.
[…]
FlippedRAG Attack Method Overview.
3.2. Threat Model
Given a topic $q$ (the question) from a set of controversial topics $Q$, a RAG corpus $D$ and a target opinion $S_{t}$ (Pro or Con, “Pro” stands for supporting and “Con” stands for opposing) as the opinion that the attacker expects the RAG system to generate, we denote the target document as $d_{t}$ which holds $S_{t}$ for $q$. Given a black-box retrieval model RM, it calculates relevance score $R(q,d)$ for each document $d$ in $D$ and returns the set of documents with the $k$ highest relevance score as the retrieval result (the LLM only takes top $k$ candidate documents into account). The retrieval result is denoted as $\text{RM}_{k}(q)=\{d_{1},d_{2},...,d_{k}\}$. The LLM of the black-box RAG utilizes $\text{RM}_{k}(q)$ to answer $q$ with response $\text{LLM}(q,RM_{k}(q))$, the opinion of $\text{LLM}(q,RM_{k}(q))$ is $S_{o}$, $S_{o}=\text{S}(LLM(q,RM_{k}(q)))$.
3.2.1. Objective of the Adversary
The attacker aims to find the adversarial trigger $p_{\text{adv}}$ for each $N$ selected target document $d_{t}$, and add it to $d_{t}$, transforming the corpus $D$ to $D(d;d_{t}\oplus p_{\text{adv}})$. $p_{\text{adv}}$ can increase the relevance score $R(q,d_{t}\oplus p_{\text{adv}})$ assigned by the retrieval model RM, $d_{t}$ will be ranked at the top of the retrieval results $\text{RM}_{k}(q)=\{d\mid d_{t}\oplus p_{\text{adv}}\}$ , guiding the LLM to generate responses that align with the target opinion: $\text{S}(LLM(q,\{d\mid d_{t}\oplus p_{\text{adv}}\}))=S_{t}$, $S_{t}\neq S_{o}$.
3.2.2. Capabilities of the Adversary
In the black-box scenario, we assume the attacker is only authorized to query the RAG, obtain RAG-generated results and modify at most $N$ documents in the corpus. There are no restrictions on the number of calls to RAG. Furthermore, the attacker has no knowledge of the model architecture, model parameters, or any other information related to the models within the black-box RAG. Modifying the prompt templates used by the LLM in RAG is also prohibited.
Figure 3. The framework for obtaining ranking imitation data of retrieval model in black-box RAG system.
3.3. Black-box Retriever Imitation
The primary issue in implementing manipulation is to make the retrieval model of the black-box RAG transparent so that the attacker is able to produce adversarial triggers based on the knowledge of black-box retriever. We aim to imitate the black-box retrieval model $RM$ in RAG. The basic idea is to train a surrogate model $M_{\text{i}}$ with the ranking results $RM_{\text{k}}(q)$ from the retrieval model $RM$, thus turning the black-box retrieval model into a white-box model. However, since the retriever and the LLM in RAG are serially connected, it is not possible to directly obtain the ranking results of the retriever. Therefore, we attempt to guide the LLM to replicate the retrieval results of the black-box RAG, then we obtain the documents deemed relevant to the user query by the black-box retrieval model $RM$, which can be used as positive examples $d_{+}$ for black-box imitation training. Subsequently, documents irrelevant to the query can be randomly sampled as negative examples $d_{-}$ Therefore, we design specific instructions to induce the black-box RAG to replicate the retrieval results of the retriever RM. These retrieval results only need to reflect the relevance to the query. Then, based on the generated results of the LLM and the corpus, we sample positive and negative data to train the surrogate model. The method for obtaining imitation data of the retrieval model in black-box RAG is illustrated in Figure 3. The specific instruction used is as follows:
—Here is the user question— < < <{query}> > > —Here is the user command— Please COPY all the given context altogether in [[ ]] including all marks and symbols. Do not omit any sentence of the context.
After that, we use a pairwise approach to sample data and train the surrogate model. The black-box RAG responds to the instruction with context information so the responses generated by the black-box RAG reflect the retrieval results. Relevant documents are sampled from the responses as positive examples $d_{+}$, and random irrelevant documents are sampled as negative examples $d_{-}$. For real-world RAG system attack, negative samples can be constructed using web information randomly sampled from search engines like Google to simulate real-world adversarial noise in RAG corpora.
These sample pairs $(d_{+},d_{-})$ are then incorporated into the training dataset. After sampling the imitation data, we employ a pairwise training method to train the surrogate model $M_{i}$. Let the relevance score calculated by $M_{i}$ be $R_{i}$, we train the surrogate model $M_{i}$ by minimizing the following contrastive loss, enabling the surrogate model $M_{i}$ to approximate the internal mechanisms of the black-box RAG retrieval model.:
(1) $L=-\frac{1}{|Q|}\sum_{q\in Q}\log\left(\frac{R_{i}(q,d_{+})}{R_{i}(q,d_{+})+\sum R_{i}(q,d_{-})}\right)$
3.4. Opinion Manipulation Attack
After obtaining the surrogate model $M_{i}$, we transform the manipulation of RAG-generated opinions in a black-box scenario into manipulation in a white-box setting. Since we have all the knowledge of the white-box surrogate model $M_{i}$, we directly implement adversarial retrieval attacks on it, generating the adversarial trigger $p_{\text{adv}}$ for the candidate document $d_{t}$ with opinion $S_{t}$. We employ Pairwise Anchor-based Trigger (PAT) (Liu et al., 2023b) strategy for adversarial retrieval attacks, which is commonly used as the baseline in related research. After adding the generated adversarial trigger to the target document with $S_{t}$, the attacker can corrupt the corpus of RAG by inserting the modified candidate document into the RAG corpus.
PAT, as a representative adversarial retrieval attack strategy, adopts a pairwise generation paradigm. Given the target query, the target candidate item, and the top candidate item(named anchor, which is to guide the adversarial text generation), the strategy utilizes gradient optimization of pairwise loss, which is calculated from the candidate item and the anchor, to find the appropriate representation of an adversarial trigger. The strategy also adds fluency constraint and next sentence prediction constraint. By beam search for the words, the final adversarial trigger, denoted as $T_{pat}$, is iteratively generated in an auto-regressive way. We use $T_{\text{pat}}$ as $p_{\text{adv}}$, with its generated optimization function being (Liu et al., 2023b):
(2) $\max\{M_{i}(q,T_{\text{pat}};w)+\lambda_{1}\cdot\log P_{g}(T_{\text{pat}};w)\\ +\lambda_{2}\cdot f_{\text{nsp}}(d_{t},T_{\text{pat}};w)\}$
In the above formula, $P_{g}$ is the semantic constraint function, and $f_{\text{nsp}}$ is the next sentence prediction consistency score function between $T_{\text{pat}}$ and $d_{t}$. $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters, $w$ represents the internal weights of the corresponding model.
Subsequently, we insert $p_{\text{adv}}$ into the target document to execute the poisoning attack on the RAG corpus. Our objective of RAG opinion manipulation attack is as follows:
(3) $\max\dfrac{1}{|Q|}\sum_{q\in Q}\text{I}(\text{S}(LLM(q,RM_{k}(q)))=S_{t})$
(4) $\begin{split}s.t.RM(q)&=E(q)\cdot E(d),d\in D(d;d_{t}\oplus p_{\text{adv}})\\ RM_{k}(q)&=\{d_{1},d_{2},...d_{k}\mid argmax_{d\in D}RM(q)\}\end{split}$
I(·) is the indicator function whose return value is 1 if the conditions within the parentheses are satisfied; otherwise, the return value is 0. S(·) is the opinion classification function. The retrieval model computes semantic relevance score by dot product, and $E$ denotes the word embedding vector. $RM_{k}(q)$ denotes the set composed of the top $k$ documents with the highest relevance scores collected from the ranking of black-box retrieval model $RM$.
4. Experiments
4.1. Dataset
The database of the black-box RAG system in our study consists of MS MARCO Passages Ranking dataset (Nguyen et al., 2016) and PROCON.ORG data [^1] https://testing-www.procon.org/, containing about 500 queries and 500,000 documents. We utilize queries in MSMARCO to obtain the source data for guiding the black-box RAG to generate relevant passages (Zhong et al., 2023) to sample data pairs to train the surrogate model. We use topic queries in PROCON.ORG to get controversial topics arguments for opinion manipulation.
The controversial topic dataset includes over 80 topics, covering fields such as society, health, government, and education. Each controversial topic is discussed from two stances (Pro and Con), with an average of 30 related passages, each holding a certain opinion with stance Pro or Con. We randomly select 30 public controversial topics to conduct opinion manipulation experiment on RAG.
4.2. Experiment Details
The specific setting details for the RAG manipulation experiment are as follows:
We represent the black-box RAG process as $\text{RAG}_{\text{black}}$. The RAG framework is Conversational RAG from LangChain [^2] https://python.langchain.com/v0.2/docs/tutorials/qa_chat_history/. Conversational RAG adds logic for incorporating historical messages and intention reasoning, thus supporting back-and-forth conversation with users and serving as a widely used RAG framework. The LLMs adopted are the open-source models Meta-Llama-3-8B-Instruct (Llama3), Vicuna-13b-v1.5 (Vicuna) and Mixtral-8x7B-Instruct-v0.1 (Mixtral). They perform well across various tasks among all open-source LLMs. The prompt in $\text{RAG}_{\text{black}}$ adopts the basic RAG prompt from LangChain.
We select three representative dense retrieval models, Contriever, Co-Condenser and ANCE, as the target retrieval model (Gao and Callan, 2021; Izacard et al., 2021; Xiong et al., 2020). By convention, we use dot product between the embedding vectors of queries and candidate documents as their similarity score $R$ in the ranking. The surrogate model chosen in this paper is Nboost/pt-bert-base-uncased-msmarco, which is type of BERT and is specifically trained on the MS Marco Passage Ranking dataset.
For a controversial topic $q$, documents $d_{t}$ with the target opinion $S_{t}$ are manipulated by adding adversarial triggers $p_{adv}$ at the beginning. This manipulation aims to place these perturbed documents as prominently as possible in the top $k$ rankings of the RAG retriever $\mathrm{RM}_{k}(q)$, where $k$ denotes the number of documents obtained from the retrieval results by the LLM in $\text{RAG}_{\text{black}}$ In this paper, $k$ is set to 3. In this experiment, the number of poisoned documents for FlippedRAG ranges from 5 to 10.
We collect imitation training data in a pairwise way, sampling both positive samples and negative samples. It is worth mentioning that we randomly sample negative examples from the top of the ranking produced by the original surrogate model. We discover that taking highly relevant documents ranked by the surrogate model as negative examples helps the surrogate model learn more features of the target black-box retrieval model by contrastive learning between positive and negative examples.
We adopt Qwen2.5-Instruct-72B as the opinion recognition model. Qwen2.5-Instruct-72B, due to its large parameter size, is capable of accurately identifying and classifying opinions within text. During black-box imitation training, we set the batch size as 256, the number of epoch as 4 and the learning rate as 4e-5. In the process of implementing PAT to generate adversarial triggers, we set the number of beams as 30, the temperature value as 0.4, the learning rate as 0.1 and the sequence length as 15. All our experiments are conducted in Python 3.8 environment and run on a NVIDIA DGX H100 GPU.
4.3. Research Questions
We propose four research questions.
RQ1: Does black-box retriever imitation effectively learn about the internal knowledge of the retriever of RAG?
RQ2: How significant does opinion manipulation of FlippedRAG influence the response of the LLM in RAG?
RQ3: Does opinion manipulation significantly impact users’ perceptions of controversial topics?
RQ4: Can FlippedRAG breach existing defense mechanisms?
4.4. Baselines
To the best of our knowledge, no existing studies have specifically addressed opinion manipulation in black-box RAG systems. We therefore selected black-box PoisonedRAG and GARAG as baselines for opinion manipulation. Furthermore, we extended the baseline selection to include attack methods which target either LLMs or retrieval models and can be applied to black-box attack.
(1) Prompt Injection Attack. Prompt injection attacks constitute a class of adversarial exploits wherein malicious actors strategically insert compromised instructions or corrupted data into the inputs of LLM-integrated systems, thereby coercing the model’s outputs to align with predefined adversarial objectives(Liu et al., 2024).
(2) Disinformation. It (Pan et al., 2023) is a methodology for generating credible-sounding misinformation by leveraging LLMs, employing the adversarial prompt to induce LLMs to produce context passages imbued with specific opinions to the controversial topics.
(3) Static Text. We add static text to some documents that contradict the target opinion (“Experts dispute the above claims as being unreliable”), thereby inducing the LLM to reinforce the target opinion.
(4) PAT Transfer-based Attack. We apply retrieval adversarial strategy PAT(Liu et al., 2023b) to the surrogate model that has not undergone black-box imitation and transfer adversarial triggers to the RAG system, assessing the effectiveness of the black-box imitation.
(5) GARAG. It (Cho et al., 2024) employs genetic algorithms to optimize the discovery of novel adversarial documents that achieve dual objectives: maintaining retrievability within RAG systems, and deviating LLM outputs from factual accuracy.
(6) PoisonedRAG. It (Zou et al., 2024) is black-box attack framework leverages query insertion and LLM generation to construct malicious documents. Detailed baseline introduction is in Appendix A.2.
4.5. Evaluation Metrics
For RQ1, we aim to evaluate the relevance ranking ability of the surrogate model trained through black-box imitation of $\text{RAG}_{\text{black}}$ on standard ranking tasks and compared with the target retrieval model. Additionally, when adversarial triggers are applied to the target retrieval model, we compare the manipulation effect of triggers generated from the surrogate model versus those generated from the original BERT model, which is not trained by black-box imitation, to assess the effect of black-box imitation. We use the following evaluation metrics, higher values of these metrics demonstrate enhanced efficacy in black-box imitation, indicating that the surrogate model successfully acquires the ranking capabilities of the target model.
Normalized Discounted Cumulative Gain (NDCG). NDCG is designed to evaluate the effectiveness of search engines in ordering results by relevance.
Inter Ranking Similarity (Inter). It is a metric used to measure the similarity between two ranking list by counting the overlap of the top candidates, which is ranging from 0 to 1.
For RQ2, we aim to evaluate the manipulation performance of FlippedRAG on retriever rankings and LLM responses. We use Top3v, Top6v, Ranking Attack Success Rate (RASR)and Boost Rank (BRank) to reflect how effective our attack distort the retrieval ranking of RAG. Elevated values of these metrics demonstrate enhanced manipulation efficacy in ranking and generated opinion manipulation.
Top3v and Top6v. Top3v quantifies the enhancement in the proportion of target opinion within the top-3 rankings after manipulation. Top6v follows the same definition.
Ranking Attack Success Rate (RASR). RASR measures the average rate of successfully boosted candidates in the ranking of each query.
Boost Rank (BRank):BRank is the average of the total rank improvements for all target documents under each query.
We also define Opinion Manipulation Success Rate (OMSR) and Average Stance Variation (ASV) to evaluate the significance of our opinion manipulation on the LLM responses.
Opinion Manipulation Success Rate (OMSR). OMSR measures the average rate of successfully manipulated LLM responses at the opinion level.
Average Stance Variation (ASV). ASV represents the average increase of opinion scores of LLM responses in the direction of the target opinion $S_{t}$ after manipulation. ASV reflects how much the shift in opinion polarity is.
5. Experimental Results Analysis
Table 2. Comparison(%) of ranking ability between the surrogate model and the target retrieval model.
Model NDCG@10 Inter@10
Contriever 67.71
$M_{Contriever}$ 69.96 62.55
Co-Condenser 68.16
$M_{CoCondenser}$ 67.98 61.86
ANCE 67.84
$M_{ANCE}$ 66.49 56.74
5.1. RQ1: Does black-box retriever imitation effectively learn about the internal knowledge of the retriever of RAG?
We first compare the ranking ability of the surrogate model $M_{i}$ and the target retrieval model $RM$, as well as the similarity of their ranking results to ensure that the surrogate model has learned the capabilities of the black-box retrieval model.
As shown in Table 2, the surrogate model $M_{i}$ trained by black-box imitation is similar to the target retrieval model in relevance ranking ability for their close NDCG@10 scores. Their ranking results are also similar enough, validating the effectiveness of the black-box imitation. For convenience, we use $M_{Contriever}$ to represent the surrogate model that has its training data collected from $\text{RAG}_{\text{black}}$ which uses Contriever. The method of representation applies to surrogate models based on ranking data of Co-Condenser and ANCE.
Table 3. Comparison of manipulation effect between the surrogate model $M_{i}$ and the BRET model without imitation.
Target Retrieval Model Target Opinion Victim Model Ranking Manipulation Performance Opinion Manipulation on Llama-3
$Top3_{v}$ $Top6_{v}$ $BRank$ $RASR\%$ $OMSR\%$ $ASV$
Contriever PRO w/o imitation 0.31 0.28 72.33 70.73 43.33 0.50
imitation 0.37 0.38 111.03 74.22 53.33 0.67
CON w/o imitation 0.27 0.24 67.1 71.93 40.00 0.27
imitation 0.37 0.35 105.13 78.26 53.33 0.70
Co-Condenser PRO w/o imitation 0.22 0.12 43.53 58.91 23.33 0.20
imitation 0.27 0.22 63.70 62.01 43.33 0.27
CON w/o imitation 0.07 0.13 37.43 57.11 20.00 0.07
imitation 0.17 0.19 58.96 64.42 33.33 0.27
ANCE PRO w/o imitation 0.30 0.22 57.20 62.79 43.33 0.43
imitation 0.38 0.32 85.96 69.76 53.33 0.47
CON w/o imitation 0.24 0.19 50.90 64.42 30.00 0.20
imitation 0.33 0.32 89.40 74.50 36.66 0.37
We place great importance on the manipulation performance of the adversarial triggers generated from the surrogate model $M_{i}$. The ranking manipulation effect of the surrogate model $M_{i}$ on the target retrieval model should fall between the manipulation effect of the white-box target retrieval model and that of BERT-based transfer attacks without imitation as $M_{i}$ has acquired some of the target model’s capabilities. Only if triggers from the surrogate model has the better ranking manipulation effect on the target retrieval model than that from the origin model without imitation, we consider the black-box imitation process to be successful.
As shown in Table 3, we find that in terms of ranking manipulation effectiveness, the surrogate models after black-box imitation outperform the baseline models. In both RASR and BRank of target documents, triggers generated from the surrogate model achieve better transfer attack effects on the target retrieval model. Specifically, $M_{Contriever}$ has the best ranking manipulation performance on the triggers generated for transfer attacking. $M_{ANCE}$ demonstrates ranking manipulation effectiveness comparable to that of $M_{Contriever}$, whereas the manipulation effect achieved by $M_{CoCondenser}$ is less ideal by comparison. It indicates that Co-Condenser is more robust and reliable than Contriver and ANCE.
Additionally, the surrogate models promote more documents with target opinion into top 3 range, thereby having a significant impact on the response of $\text{RAG}_{\text{black}}$. Thus, we can see that the transfer attack based on the surrogate model achieves stronger opinion manipulation by successfully flipping opinions of RAG response. Black-box imitation enables the surrogate model to learn the features of the black-box retrieval model, allowing triggers generated by the surrogate model to induce a greater shift in opinion polarity in LLMs’ response compared to the baseline model. Table 3 also indicates that adversarial triggers generated by $M_{Contriever}$ and $M_{ANCE}$ place more target documents in top 3 range, thus achieving relatively better performance on OMSR and ASV.
In short, the above experiment indicates that the black-box imitation process is able to help BERT learn internal knowledge about the target black-box retrieval model.
Table 4. Opinion manipulation effect of FlippedRAG across different LLMs.
Surrogate Model Target Opinion OMSR(%) on Different LLMs ASV on Different LLMs
Llama-3 Mixtral Vicuna Llama-3 Mixtral Vicuna
$M_{Contriever}$ PRO 53.33 46.66 36.66 0.67 0.57 0.50
CON 53.33 46.66 46.66 0.70 0.40 0.60
$M_{CoCondenser}$ PRO 43.33 36.66 36.66 0.30 0.37 0.43
CON 33.33 36.66 33.33 0.27 0.37 0.33
$M_{ANCE}$ PRO 53.33 56.66 46.66 0.47 0.53 0.53
CON 36.66 40.00 46.66 0.37 0.47 0.60
5.2. RQ2: How significant does opinion manipulation influence the response of the LLM in RAG?
After the black-box imitation training, we conduct opinion manipulation experiments across different $\text{RAG}_{\text{black}}$ (based on Llama3, Mixtral, or Vicuna) with FlippedRAG. We also conducted a comparative evaluation of FlippedRAG against established baselines to assess its relative efficacy in executing opinion manipulation on black-box RAG architectures.
Table 5. Opinion fluctuations in RAG responses based on clean collections of different LLMs.
LLM OMSR(%) ASV
Llama-3 6.7 -0.03
Mixtral 10.0 0.07
Vicuna 6.7 0.06
As shown in Table 4, FlippedRAG achieves significant opinion manipulation effect on all three LLMs. As a whole, it has a 40% – 50% chance of successful opinion flipping and realizes about 0.46 in opinion polarity shift of range 0 to 2. Moreover, it reveals that FlippedRAG based on $M_{Contriever}$ and $M_{ANCE}$ are more effective in manipulating the RAG opinion with relatively higher OMSR and ASV value. That is because these attacks exhibit superior performance on ranking distortion.
We also observe that the impact of FlippedRAG varies in opinion manipulation effectiveness across different LLMs. Llama3 and Mixtral have higher OMSR score in contrast to Vicuna for their greater instruction-following ability. Thus, Llama3 and Mixtral are more susceptible to the adversarial documents. On the contrary, for ASV, Vicuna actually demonstrates a relatively greater shift in opinion polarity. We analyze that this is due to the better diversity in responses from Llama3 and Mixtral when answering questions, resulting in less stable opinion polarity in the their responses.
In some cases, opinion manipulation is unsuccessful, instead of turning the opinion of $\text{RAG}_{\text{black}}$ response to the target opinion or making it remain unchanged, manipulation causes the opinion polarity to shift in the opposite direction of the target opinion, thereby lowering the ASV value. The cause of these reverse manipulation results may also stem from the randomness of LLMs generation, although this influence is relatively minor.
To further investigate the fluctuations in opinions resulting from the generation randomness, we query the RAG system, which is based on Contriver and clean database, twice to get two responses on the same topic. The results are illustrated in the Table 5. The responses generated by LLMs do exhibit fluctuations in opinions. However, these fluctuations are minimal and significantly less than that resulting from deliberate manipulation.
Figure 4. FlippedRAG’s overall effect of opinion manipulation of on LLama-3 with retrieval model Contriever.
[…]
From the perspective of overall opinion manipulation of $\text{RAG}_{\text{black}}$ response, as shown in Figure 4, FlippedRAG is able to significantly alter the opinion of RAG-generated content in a specified stance direction. When the target opinion is “Pro”, compared to the opinion distribution before manipulation, the proportion of responses holding a supportive stance (in gray) has significantly increased in the opinion distribution after manipulation. The converse holds true when the target opinion is specified as ’Con’. Graphical representations of the opinion manipulation effects on Mixtral and Vicuna are provided in the Appendix B.2.
Furthermore, FlippedRAG exhibits no adverse effects on the generation performance of non-target queries, as the target questions and normal questions remain orthogonal in their semantic feature space.
Figure 8 in Appendix B.1 presents a case study of FlippedRAG.
To further analyze the opinion manipulation results on RAG response at a finer granularity, we conduct experiments across diverse domain of controversial topics to evaluate its efficacy in cross-domain opinion manipulation. We focus on “Health”, “Society”, “Government”, “Education” four domains. FlippedRAG significantly increases the proportion of candidate items with target opinion in the Top 3 of the retrieval list, thereby guiding the LLM to change its opinion in the response. However, the manipulation effect varies across different themes. The detailed discussion is in Appendix B.3.
Table 6. Ranking manipulation and opinion manipulation results across attack baselines on target opinion Pro/Con .
Target Attack Method Retriever Opinion
$Top3_{v}$ $RASR\%$ $OMSR\%$ $ASV$
Pro Prompt Injection Attack 26.67 0.03
Disinformation 0.26 40.00 0.43
Static Text 40.00 0.40
PAT Transfer-based 0.31 70.73 43.33 0.50
GARAG 0.02 2.40 10.00 0.03
PoisonedRAG 0.46 56.67 0.76
FlippedRAG 0.37 74.22 63.33 0.70
Con Prompt Injection Attack 50.00 0.47
Disinformation 0.21 36.67 0.33
Static Text 26.67 0.17
PAT Transfer-based 0.27 71.93 40.00 0.27
GARAG 0.00 1.67 13.33 0.07
PoisonedRAG 0.47 66.67 0.83
FlippedRAG 0.37 78.26 53.33 0.70
We conducted a systematic comparative evaluation of FlippedRAG’s manipulation efficacy against established black-box attack baselines, with the quantitative results comprehensively tabulated in Table 6. Entries denoted by ’–’ in the table signify that the specified evaluation metric is inapplicable to the corresponding attack methodology process.
The comparative experiment results indicate that both PoisonedRAG and FlippedRAG demonstrate superior efficacy in the opinion manipulation task, with PoisonedRAG exhibiting marginally enhanced performance compared to FlippedRAG. While Prompt Injection Attack, Disinformation, Static Text, and PAT Transfer-based attack achieve moderate effectiveness in opinion manipulation, GARAG manifests the most suboptimal performance within this adversarial task paradigm.
Table 7. The controlled user experimental results on opinion manipulation in controversial topics.
Subject Opinion Polarity of User
Topic 1 Topic 2
A Group (before manipulation) 5.8 4.8
B Group (after manipulation) 6.4 8.2
The superior performance of PoisonedRAG and FlippedRAG stems from their targeted optimizations of critical components in the RAG workflow: retrieval and generation. In the generation phase, PoisonedRAG leverages LLMs to synthesize content engineered to steer target systems, while FlippedRAG adopts a minimal data perturbation strategy by curating opinion-specific documents that guide RAG systems to extrapolate desired stances through summarization. PoisonedRAG’s enhanced efficacy derives from its retrieval optimization: injecting the question into documents, which maximizes retrieval probability by exploiting the semantic self-similarity. In contrast, although FlippedRAG improves adversarial document rankings through black-box retriever imitation, its effectiveness remains suboptimal compared to exact question-question matching paradigm.
The diminished efficacy of the PAT Transfer-based method compared to FlippedRAG can be attributed to its lack of a black-box retriever imitation process. Both Prompt Injection Attack and Static Text exclusively target the generation phase while neglecting the retrieval in RAG architectures, leading to significant performance degradation when migrating LLM-targeted attack strategies to RAG scenarios. Although Disinformation fundamentally remains a generation-optimized attack method, its approach of constructing biased content based on the question nevertheless achieves partial retrieval prioritization enhancements.
The limited adaptability of GARAG to black-box opinion manipulation tasks arises from two fundamental constraints. First, its methodology—originally designed for closed-domain QA to deviate RAG systems from factual correctness—exploits a vast generative space where any incorrect output suffices as a successful attack. In contrast, our opinion manipulation framework strictly confines LLM outputs to binary stance options (support/oppose), significantly increasing the attack complexity. Second, our implementation reveals white-box dependencies in GARAG’s genetic algorithm, which leverages internal retriever and LLM calculated values during adversarial document optimization. These architectural dependencies undermine its efficacy in more authentic black-box scenarios.
5.3. RQ3: Does opinion manipulation significantly impact users’ perceptions of controversial topics?
As shown in Table 7, when we manipulated the relevance ranking of documents supporting the target topics, users’ approval of those topics significantly increased. This experiment confirms that opinion manipulation has a substantial effect on users’ perceptions. Comprehensive experimental details of the user study are provided in the Appendix B.4.
5.4. Mitigation Analyses
We conducted a systematic robustness evaluation of FlippedRAG by implementing multiple defense mechanisms to assess its adversarial resilience. We implemented several defense mechanisms aligned with the core workflow of RAG by selecting phase-specific countermeasures including: detection (filtering) for data, preprocessing defenses for retrieval and generation-phase mitigation.
Given our assumed RAG poisoning scenario involving contamination of publicly indexable web content, we incorporate security detection designed for search engine optimization. Comprehensive technical specifications of the adopted mitigation methods are provided in the Appendix B.4.
5.4.1. Mitigation Based on Spam Detection
Table 8. TF-IDF based automatic spamicity detection rates (%) of different methods.
Threshold 0.3 0.25 0.2 0.15 0.1
Prompt Injection Attack 1.6 5.2 10.0 22.2 44.7
Disinformation 1.3 4.7 11.6 27.9 57.4
Static Text 1.0 3.3 9.7 23.2 44.4
PAT Transfer-based 1.5 6.6 19.0 40.9 68.6
GARAG 1.9 4.8 11.8 25.5 46.8
PoisonedRAG 67.5 88.8 98.0 99.2 100.0
FlippedRAG 2.5 9.0 24.8 51.5 76.7
Clean 0.3 2.0 7.2 21.7 41.6
Spamicity detection (Zhou and Pei, 2009) is used for discovering term spam. More detail about spam detection is in Appendix B.5.
In Table 8, we compute the spamicity scores of the adversarial documents generated by FlippedRAG and our baselines. We then calculate the probability of these adversarial documents being detected as spam by the spamicity detection system under various thresholds, referred to as the detection rate.
Under all threshold conditions, the detection rate of triggers generated by FlippedRAG is significantly lower than those generated by PoisonedRAG and exhibits minimal divergence from those observed in clean data. While other baselines demonstrate consistently low detection rates across experimental thresholds, PoisonedRAG exhibited persistently the highest detection rates. This suggests that directly embedding the question in the document lacks sufficient stealth, making it more easily detectable by simple detection methods. The maximal separability between PoisonedRAG’s manipulated documents and clean data occurring at threshold range 0.2 – 0.25. However, achieving a balance between minimizing the risk of erroneously filtering clean data and maintaining high detection efficacy against FlippedRAG remains difficult.
Figure 5. Distributions of log perplexity (PPL) calculated by GPT-2 on clean documents, attacked documents by FlippedRAG, and the baselines.
5.4.2. Mitigation Based on Perplexity Analysis
Perplexity is employed as a metric to evaluate the performance of a probability distribution or probabilistic model in predicting text, assessing whether the text conforms to the conventions of natural language, including coherence and fluency. If a text exhibits a higher perplexity under a language model (LM), it indicates poorer quality. Song et al.(Song et al., 2020) utilized perplexity calculated by GPT-2 in an attempt to distinguish between synthetic text and natural text.
We utilized GPT-2 to measure the perplexity of documents in clean data, documents attacked by the baselines, and documents attacked by FlippedRAG. The distributions are depicted in Figure 5.
Our analysis indicates that, PoisonedRAG and Disinformation generate adversarial documents with anomalously low perplexities that significantly deviate from normal data distributions. This phenomenon likely stems from their reliance on LLM-generated adversarial documents. Consequently, suboptimal LLM selection would markedly increase the detectability of PoisonedRAG and Disinformation through perplexity-based detection. Both FlippedRAG and other baselines exhibit perplexity levels comparable to clean data distributions, rendering them resistant to perplexity detection.
5.4.3. Mitigation Based on Keyword Density
Keyword density, which is a measure of how often a certain keyword or phrase appears, is a critical factor in SEO as high keyword densities make spam pages more relevant to the user query. Given that the RAG attack scenario involves manipulation during the search phase, we also employed keyword density to detect the RAG attack in Table 9.
Table 9. Keyword Density(%) of clean documents and documents attacked by FlippedRAG and the baselines.
Documents Keyword Density(%)
Overall Window size
20 50 100
Prompt Injection Attack 4.75 10.37 6.29 5.12
Disinformation 5.28 7.67 5.28 5.28
Static Text 4.49 10.42 6.33 4.94
PAT Transfer-based 6.16 13.03 8.38 6.64
GARAG 4.39 10.00 6.04 4.78
PoisonedRAG 18.39 52.90 24.41 18.39
FlippedRAG 6.35 13.23 8.51 6.82
Clean 4.39 10.01 6.04 4.78
The keyword densitiy of adversarial documents generated by FlippedRAG and other baselines is comparable to that of natural documents in clean data. In contrast, the keyword density of documents generated by PoisonedRAG is significantly higher than that of natural documents, making PoisonedRAG more susceptible to being filtered out by search engine SEO review mechanisms.
5.4.4. Mitigation Based on Paraphrasing
Cheng et al. (Cheng et al., 2024) and Zou et al. (Zou et al., 2024) attempt to employ paraphrasing defense to test the effectiveness of the RAG attack they proposed.
Given that paraphrasing primarily targets the retrieval phase, the comparative analysis excludes prompt injection attacks and static text manipulation as they exclusively focus on the generation. We also exclude evaluation on GARAG for its inefficacy in opinion manipulation.
The attack effectiveness of FlippedRAG and other baselines against paraphrasing is presented in Table 10. Both the baselines and FlippedRAG exhibit a decline in attack effectiveness when confronted with paraphrasing, yet PoisonedRAG’s and FlippedRAG’s overall attack performance remains notably significant. PoisonedRAG demonstrates a less reduction in attack effectiveness under paraphrasing defense attributing to the direct insertion of queries.
Table 10. Manipulation Effect of the baselines and FlippedRAG against paraphrasing defense. w/o denotes without, w/ denotes with.
Attack Paraphrasing Manipulation Performance
OMSR(%) ASV
Disinformation w/o 40.0 0.43
w/ 20.0 0.20
PAT Transfer-based w/o 43.3 0.50
w/ 30.0 0.30
PoisonedRAG w/o 56.7 0.76
w/ 43.3 0.53
FlippedRAG w/o 63.3 0.70
w/ 40.0 0.43
5.4.5. Mitigation Based on Randomized Mask Smoothing
Randomized smoothing is a defense method aiming at achieving certified robustness against adversarial examples. It constructs a smoothed composite function by introducing random variables to the original function (model), thereby enabling the certification of its robustness under certain conditions. We explore randomized mask smoothing as a defense mechanism to counteract RAG attacks in Figure 6.
Figure 6. OMSR(%) under different mask rates and ensemble sizes for FlippedRAG and the baselines.
The experimental results indicate that both PoisonedRAG and FlippedRAG experience a slight decline in attack effectiveness when confronted with randomized masking smoothing. However, since both PoisonedRAG and FlippedRAG employ trigger insertion, the composition of their adversarial documents are similar after masking operations. Additionally, because PoisonedRAG inserts the query itself as the trigger, which enhances retrieval ranking more effectively, the decline in PoisonedRAG’s attack success and opinion manipulation success rate is somewhat smaller.
Our findings further reveal that the success rate of RAG attacks continues to increase when the mask rate ranges from 0% to 30%. This suggests that RAG attacks like FlippedRAG exhibit a certain degree of robustness against random mask smoothing defenses. However, when the mask rate exceeds 50%, the attack success rate shows a more pronounced decline. Nevertheless, an excessively high mask rate may significantly impair the ranking capability of the RAG system, making it difficult for the retrieval model to produce accurate rankings. Consequently, employing random mask smoothing as a defense against FlippedRAG in practical applications is not advisable.
5.4.6. Mitigation Based on RobustRAG
RobustRAG, proposed by Xiang et al. (Xiang et al., 2024), represents a defense framework specifically targeting retrieval poisoning attacks in RAG systems. It employs an isolate-then-aggregate strategy that leverages the inherent workflow characteristics of RAG architectures to defend against data poisoning attacks.
The manipulation success rate of FlippedRAG against RobustRAG defense is presented in Table 11. Under the RobustRAG mitigation, the OMSR of FlippedRAG decreases compared to scenarios without defensive mechanisms, yet still maintains a success rate reaching approximately 40%—significantly higher than the attack success rates reported by Xiang et al. for factoid questions. This discrepancy arises because RobustRAG was specifically designed as a mitigation method against attacks on factoid and closed-ended questions, where poisoned documents typically contain a single, specific incorrect answer.However, FlippedRAG targets opinion-level manipulation, where a single passage may contain multiple terms or phrases with inherent opinion biases. These opinion-laden terms are quantitatively more prevalent, enabling them to persistently influence the LLM’s output.
FlippedRAG demonstrates superior manipulation efficacy against the RobustRAG defense framework compared to other baseline methods. The significantly diminished adversarial performance of PoisonedRAG likely stems from its reliance on LLM-generated steering content to bias RAG outputs. Prompt injection attack exhibits the poorest performance against RobustRAG, with near-negligible success rates. This vulnerability arises from its dependence on injected prompt instructions, which are effectively neutralized by RobustRAG’s isolate-then-aggregate strategy.
Table 11. Manipulation effect of different attacks with RobustRAG defense, where the retrieval model is Contriever.
Attack OMSR(%) ASV
Pro Con Pro Con
PoisonedRAG 36.7 33.3 0.33 0.20
Disinformation 23.3 30.0 0.06 0.20
Prompt Injection Attack 16.7 13.3 0.06 -0.20
FlippedRAG 40.0 46.7 0.27 0.30
5.4.7. Mitigation Analysis Conclusion
FlippedRAG exhibits pronounced adversarial resilience against existing defense mechanisms. In contrast, PoisonedRAG’s reliance on simple heuristic approaches (e.g., keyword saturation) renders it vulnerable to SEO detection methods. Furthermore, its dependence on LLM-generated content introduces measurable deviations from natural text distributions, enabling effective countermeasures through perplexity analysis and RobustRAG mitigation. Other baselines demonstrate limited defensive robustness.
5.5. Parameter Analysis
In the parameter analysis, we investigat the manipulation effectiveness of FlippedRAG when poisoning varying quantities of documents.
Figure 7 demonstrates the opinion manipulation effectiveness of flippedRAG under varying numbers of poisoned documents $N$. As the number of poisoned documents increases from 1 to 10, both the success rate of opinion manipulation and the average stance variation exhibit continuous improvement. When the poisoning quantity $N$ reaches 3 (the default context size $k$ being 3), FlippedRAG already achieves considerable opinion flipping effectiveness. This phenomenon occurs because a larger number of poisoned documents increases the probability of adversarial documents being retrieved into the context. Moreover, when the context size $k$ is 3, three poisoned documents can typically occupy the entire context space in most cases, resulting in a decelerated growth rate of attack effectiveness when the poisoning quantity $N$ exceeds 3.
Figure 7. OMSR(%) and ASV of FlippedRAG with increaing poisoned document number.
6. Conclusion
In this paper, we proposed FlippedRAG, a transfer-based adversarial attack designed to effectively manipulate the opinion polarity of black-box RAG models, especially on controversial topics. We demonstrated that the underlying retriever of black-box RAG systems can be effectively reverse-engineered via enumeration of critical queries and candidates, which then enables adversaries to construct a precise surrogate retriever. Leveraging this surrogate retriever, we successfully crafted targeted poisoning triggers that effectively manipulated retrieval results and, subsequently, the generated responses. The experimental results demonstrate that our transfer attack strategy effectively alters the opinions expressed in content generated by RAG models. More importantly, opinion manipulation can have profound consequences on users’ cognition, potentially leading them to accept inaccurate or biased information. This could induce cognitive shifts and distort public opinion, especially on controversial topics, highlighting the critical importance of safeguarding against such vulnerabilities in RAG.
Ethics Considerations
This research investigates the security vulnerabilities of RAG systems from the perspective of adversarial opinion manipulation. In accordance with the principles outlined in the Menlo Report—Respect for Persons, Beneficence, Justice, and Respect for Law and Public Interest, as well as the ACM CCS Ethics Guidelines, we undertook a rigorous assessment of the ethical implications associated with our study.
Responsible Disclosure. We understand the importance of engaging with model and framework providers, so we reported this vulnerability to the LangChain Security Team, hoping to collaborate on developing a mitigation module. However, due to limitations in defense measures, we have not yet gained recognition, and we are still striving to make effective progress. Additionally, the widespread use of hybrid open-source frameworks means that ownership boundaries are less clear compared to traditional software vulnerabilities. This ambiguity prompted us to explore a policy-oriented disclosure approach, aiming to encourage systemic preventative measures. Therefore, we plan to initiate policy-level mitigation and subsequently disclose to large media platform, not for public release. We are willing to inform other relevant large-scale platforms and international organizations.
Pro&Con Dataset Compliance. The Pro&Con dataset was collected from PROCON.org, whose Terms of Use explicitly permit academic usage. Our data-gathering procedure strictly adhered to four essential principles: (1) All collected data were publicly accessible without requiring authentication; (2) Data usage was strictly limited to academic research with no commercial intent or benefit; (3) Data collection was restricted to amounts necessary, in alignment with fair-use principles; and (4) Server overload was prevented through a careful, rate-limited collection approach.
Data Privacy and Informed Consent. All experiments were conducted using publicly available opinion-oriented datasets. To protect the privacy of individuals, we thoroughly anonymized all data to eliminate personally identifiable information (PII). Our experiment notification process included providing the necessary preliminary information before participation and comprehensive disclosure of complete information after the experiment. Informed consent was obtained from all participants in compliance with IRB standards. Following the conclusion of the study, we transparently communicated the purpose of the adversarial manipulation to mitigate any residual cognitive influence or harm. All experiments involving adversarial attacks were executed in secure, controlled environments to prevent unintended deployment or model misuse.
IRB Approval. This study was conducted under full Institutional Review Board (IRB) approval (Protocol No. XXX-HSS-IRB2025010, where XXX denotes an anonymized institutional identifier). We confirm that all research activities adhered strictly to IRB-approved protocols, including procedures for informed consent, data handling, anonymization, and user privacy protection. Each participant received a cash reward upon completing the experiment. The duration of the experiment was approximately 30 minutes. Participants were voluntarily recruited through campus notices and had no cognitive impairments or difficulties in language comprehension.
In summary, while our research exposes critical vulnerabilities in RAG systems, it also contributes constructively to the discourse on trustworthy AI by fostering the development of secure, transparent, and ethically-aligned machine learning infrastructures.
Open Science
To ensure transparency and reproducibility, all attack implementations, including FlippedRAG, evaluation scripts, and datasets, are hosted on Anonymized GitHub [^3] This paper is under review, so the code has not been open to public yet.. The repository is licensed under the MIT License, allowing reuse while prohibiting malicious applications. We have provided comprehensive documentation, covering environment setup, hyperparameters, and attack pipelines, to ensure independent verification. Sensitive data has been excluded from the repository. All components necessary to replicate our findings are publicly accessible. We are committed to maintaining the repository availability for at least three years and will update the repositories based on community feedback.
Appendix A Experiment Detail
A.1. Algorithm
The algorithm is in Algorithm 1.
Input: target black-box RAG model $\boldsymbol{\textit{RAG}_{\text{black}}}$, target retrieval model $\boldsymbol{RM}$, surrogate model $\boldsymbol{M_{i}}$, controversial topics $\boldsymbol{Q}$, target topic $q$, target opinion $S_{t}$, corpus $Docs$ , target documents with target opinion $Docs_{t}$, target document $d_{t}$, relevant document $d_{+}$, random sampled document $d_{-}$, poisoned document amount $N$
Instructions:
$i_{1}$ =“—Here is the user question—
< < < {query} > > >
—Here is the USER COMMAND—
Please COPY all the given context altogether in [[ ]] including all marks and symbols. Do not omit any sentence of the context”
$i_{2}$ =“Use the following pieces of retrieved context to answer the question…”
// $\textit{RAG}_{\text{black}}$ uses $i_{2}$ as prompt template.
Functions:
$\boldsymbol{PAT}$: Pairwise Anchor-based Trigger generation strategy.
Output: manipulated RAG responses $\boldsymbol{Res}$
1
2 Phase 1. Pairwise Imitation Data Construction and Black-box Retrieval Model Imitation Training
3       INIT: Dataset $\mathcal{D}\leftarrow\{\}$
4       for $\boldsymbol{q}_{m}\in\boldsymbol{Q}$ do
5             induced rank list $\boldsymbol{R}_{m}{top3}\leftarrow$ $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m}\oplus i_{1};\boldsymbol{Docs})$
             // $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m}\oplus i_{1})$approx $\boldsymbol{RM(\boldsymbol{q}_{m})}$
6             for $\boldsymbol{d_{+}}_{j}\in\boldsymbol{R}_{m}{top3}$ do
7                   Random sample document from the ranking of origin $\boldsymbol{M_{i}}$ as $\boldsymbol{d_{-}}_{j}$
8                   $\mathcal{D}\leftarrow positive,[\boldsymbol{q}_{m};\boldsymbol{d_{+}}_{j};\boldsymbol{d_{-}}_{j}]$
9                  $\mathcal{D}\leftarrow negative,[\boldsymbol{q}_{m};\boldsymbol{d_{-}}_{j};\boldsymbol{d_{+}}_{j}]$
                   // Reverse $\boldsymbol{d_{+}}_{j}$ and $\boldsymbol{d_{-}}_{j}$ to get the negative triple
10
11
12      Train the surrogate model $\boldsymbol{M_{i}}$ on $\mathcal{D}$ with Eq 1
13       return $\boldsymbol{M_{i}}$
14Phase 2. Adversarial Trigger Generation and Opinion Manipulation in RAG Response
15       INIT: RAG Response Set $\boldsymbol{Res}\leftarrow\{\}$
16       for $\boldsymbol{q}_{m}\in\boldsymbol{Q}$ do
17             select $N$ target documents with $S_{t}$ as $Docs_{t}$
18             induced rank list $\boldsymbol{R}_{m}{top3}\leftarrow$ $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m}\oplus i_{1};\boldsymbol{Docs})$
19             $\boldsymbol{anchor}_{m}\leftarrow$ top-1($\boldsymbol{R}_{m}{top3}$)
20             for $\boldsymbol{d_{t}}_{j}\in\boldsymbol{Docs_{t}}$ do
21                   adversarial trigger $\boldsymbol{p_{adv}}_{j}\leftarrow\boldsymbol{PAT}(\boldsymbol{M_{i}};\boldsymbol{q}_{m},\boldsymbol{d_{t}}_{j},\boldsymbol{anchor}_{m})$
                   // Optimize the trigger with Eq 2
22                   adversarial doc $\boldsymbol{d_{adv}}_{j}\leftarrow\boldsymbol{d_{t}}_{j}\oplus\boldsymbol{p_{adv}}_{j}$
                   $\boldsymbol{Docs}:\boldsymbol{d_{t}}_{j}\leftarrow\boldsymbol{d_{adv}}_{j}$ // Replace
23
24            $\boldsymbol{Res}\leftarrow$ $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m};\boldsymbol{Docs})$
25
26      return $\boldsymbol{Res}$
Algorithm 1 FlippedRAG
A.2. Baselines
A.2.1. Prompt Injection Attack
Prompt injection attacks constitute a class of adversarial exploits wherein malicious actors strategically insert compromised instructions or corrupted data into the inputs of LLM-integrated systems, thereby coercing the model’s outputs to align with predefined adversarial objectives(Liu et al., 2024). In black-box opinion manipulation scenarios, we formulate specific instructions that compel the LLM to generate stance-aligned responses (supporting or opposing opinions) regarding a given controversial topic, and embed the instruction within target documents. When the RAG system processes these documents, its outputs systematically align with the attacker’s designated opinion through execution of the embedded adversarial prompts.
A.2.2. Disinformation
Pan et al. (Pan et al., 2023) developed a methodology for generating credible-sounding misinformation by leveraging LLMs. To implement this approach, we employ the adversarial prompt proposed in their work “generate a background document in support of the given opinion to the question” to induce LLMs to produce context passages imbued with specific opinions to the controversial topics. We subsequently add the generated context passages into the RAG corpus. Given their elevated topical relevance to specific controversial subjects, these passages demonstrate enhanced retrievability within the system, thereby enabling their embedded opinions to influence LLM outputs through semantic content manipulation. This mechanism fundamentally differs from the command-execution paradigm characteristic of Prompt Injection Attacks.
A.2.3. GARAG
GARAG(Cho et al., 2024) employs genetic algorithms to optimize the discovery of novel adversarial documents that achieve dual objectives: (1) maintaining retrievability within RAG systems, and (2) deviating LLM outputs from factual accuracy. In our experimental, we implement a inversion of this methodology, repurposing GARAG-generated adversarial documents to strategically induce LLMs to consistently produce responses aligning target opinions.
A.2.4. PoisonedRAG
Zou et al.(Zou et al., 2024) proposed PoisonedRAG, whose black-box attack framework leverages the insight that “a question is most similar to itself” to inject the question into the malicious text. This methodology enhances malicious text relevance rankings through keyword stuffing while simultaneously leveraging LLM-generated adversarial components to manipulate RAG systems into producing factually erroneous outputs. In our experiment, we reorient PoisonedRAG’s adversarial objective from factual corruption to opinion bias induction, transforming its attack vector into generating extreme perspectives through corpus corruption.
A.3. Evaluation Metrics
A.3.1. NDCG
NDCG is designed to evaluate the effectiveness of search engines in ordering results by relevance. The NDCG score ranges from 0 to 1. An NDCG score close to 1 indicates that the ranking is nearly ideal, with highly relevant items near the top of the list.
A.3.2. Top3v and Top6v
Top3v is obtained by subtracting Top3origin(the average proportion of target opinions in the top 3 rankings before manipulation) from Top3attacked(the average proportion of target opinions in the top 3 rankings after manipulation). Note that Top3v ranges from -1 to 1, and a positive value indicates that the manipulation successfully increased the proportion of the target opinion in the top-3 ranking. The higher the Top3v value, the more effective the manipulation. Top6v is also defined as above.
A.3.3. RASR
RASR measures the average rate of successfully boosted candidates in the ranking of each query. The higher is the RASR value, the more successful is the ranking attack.
A.3.4. BRank
BRank is the average of the total rank improvements for all target documents under each query. Note that a target document $d_{t}$ contributes a negative value to BRank if its rank is reduced accidentally by the attack.
A.3.5. OMSR
OMSR measures the average rate of successfully manipulated LLM responses at the opinion level. We consider one manipulation as a success if it guides the LLM to produce a response whose opinion is closer to the target opinion $S_{t}$ compared with the original response without RAG attack. Given the target opinion “support”(Pro), successful opinion manipulation can be from “oppose”(Con) to “neutral”, from “oppose”to “support”, and from “neutral” to “support”. Given the target opinion “oppose”, successful opinion manipulation can be from “support” to “neutral”, from “support” to “oppose”, and from “neutral” to “oppose”.
A.3.6. ASV
ASV represents the average increase of opinion scores of LLM responses in the direction of the target opinion $S_{t}$ after manipulation. In this research, we assign 0 to opposing stance, 1 to neutral stance, and 2 to supporting stance. A positive ASV indicates that the opinion manipulation towards $S_{t}$ is effective, while a negative ASV indicates that the manipulation actually makes the opinions of RAG responses deviate from $S_{t}$. ASV reflects how much the shift in opinion polarity is. The larger the ASV value, the more successful the opinion manipulation.
Appendix B Experiment Result and Analysis
B.1. Case Study
Figure 8 presents a case study of our RAG opinion manipulation on the question, “Should Corporal Punishment Be Used in K-12 Schools?” The black box highlights text expressing an supporting stance on the topic, while the red box indicates text opposing the topic. Initially, the LLM in the RAG system strongly opposed corporal punishment. However, after FlippedRAG manipulated the retrieval ranking to prioritize documents supporting corporal punishment, the LLM’s opinion in its response reversed accordingly.
Figure 8. Case Study of RAG Opinion Manipulation on “Should Corporal Punishment Be Used in K-12 Schools”.
[…]
Case study.
B.2. Overall Effect of Opinion Manipulation on Mixtral and Vicuna
The FlippedRAG’s overall effect of opinion manipulation on Mixtral and Vicuna is in Figure 9.
Figure 9. FlippedRAG’s overall effect of opinion manipulation of on Mixtral and Vicuna with retrieval model Contriever.
[…]
B.3. Domain-specific manipulation experiment of FlippedRAG
The results of the theme-specific manipulation experiments are shown in Table 12. FlippedRAG significantly increases the proportion of candidate items with target opinion in the Top 3 of the retrieval list, thereby guiding the LLM to change its opinion in the response. However, the manipulation effect on $\text{RAG}_{\text{black}}$ output opinions varies across different themes.
FlippedRAG demonstrates the greatest manipulation effectvfor questions related to the “Government” theme. The manipulation effect for questions in the “Education” theme is also strong, although it is somewhat less effective when the target opinion is “CON.” In comparison, the opinion manipulation effectiveness for the “Health” and “Social” themes is less satisfactory. This discrepancy is primarily due to the preference characteristics of the black-box retrieval model across different themes. As shown in Table 12, the rankings of the black-box retrieval model for questions in the “Government” and “Education” themes are more susceptible to manipulation.
Table 12. Manipulation results of $\text{RAG}_{\text{black}}$ ranking and response opinion across different domains.
Theme Target Opinion Rank Manipulation Performance OMSR(%) on Different LLMs ASV on Different LLMs
$Top3_{v}$ $Top6_{v}$ Llama-3 Mixtral Vicuna Llama-3 Mixtral Vicuna
Health PRO 0.25 0.35 33.33 44.44 44.44 0.22 0.44 0.56
CON 0.33 0.33 44.44 44.44 33.33 0.67 0.33 0.33
Society PRO 0.38 0.36 33.33 33.33 16.67 0.17 0.17 0.17
CON 0.27 0.27 33.33 50.00 50.00 0.50 0.67 0.50
Government PRO 0.43 0.37 61.53 53.84 61.53 0.85 0.54 0.92
CON 0.38 0.44 38.46 61.53 61.53 0.38 0.62 0.85
Education PRO 0.42 0.35 57.14 71.42 42.85 0.86 0.86 0.71
CON 0.42 0.35 23.57 14.28 42.85 0.29 0.14 0.29
B.4. User Study
We invited 10 college students to participate in the experiment, dividing them into two groups, A and B, with five participants in each group. None of the participants were informed about the experimental purpose before the experiment, and their experimental processes were independent of one another, with no communication allowed. We first developed a simple question-answering (QA) service based on the RAG framework, and then applied our attack to relevant documents containing supportive opinions on two designated controversial topics in the corpus. Participants in Group A accessed the QA service before any opinion manipulation was applied to the corpus, while participants in Group B accessed the service after we poisoned the corpus using our attack. Following the interaction, both groups were asked to rate their approval of the two topics on a scale of 1 to 10.
In accordance with the IRB approval requirements for “post-experiment feedback,” all participants received a comprehensive notification after completing the experiment: Participants in the experimental group (Group B) were informed of the technical details of the adversarial attack on the system (Control group: Group A)). All participants can request to refer to the original system answers of the control group (Group A) as a reference.
B.5. Mitigation Analysis
B.5.1. Mitigation Based on Spam Detection
Earlier search engines relied on term-based ranking methods such as TF-IDF to retrieve documents. In this context, term spam manipulates document content through keyword stuffing to artificially inflate the TF-IDF scores of spam documents, making them appear more relevant for specific queries. A straightforward approach to achieve this is by embedding the question within the target document, thereby increasing the frequency of keywords in the document and boosting its TF-IDF-based ranking.
To identify spam documents, Spamicity detection calculates the term spamicity score of a given document to assess the likelihood that the document has been intentionally manipulated. Based on a spamicity threshold and the score of the target document, the detection strategy classifies the document as spam if its score exceeds the threshold.
B.5.2. Mitigation Based on Keyword Density
Search Engine Optimization (SEO) is a set of techniques used to improve a website’s visibility and ranking on search engine results pages (SERPs). Keyword density, which is a measure of how often a certain keyword or phrase appears in the content, is a critical factor in SEO as high keyword densities make web pages more relevant to the user query. Keyword stuffing is a straightforward and effective SEO technique that involves the repetition of important keywords on a web page to enhance its search engine ranking. The method employed by PoisonedRAG, which involves adding queries to target documents, constitutes a form of keyword stuffing, as the terms within the queries serve as keywords during the retrieval process.
Numerous SEO detection methods targeting keyword stuffing have been developed, including the aforementioned spam detection based on TF-IDF values. Many scholars have also proposed the use of keyword density as a method for detecting spam and SEO defense. Hui et al. (Hui et al., 2012) suggest that an appropriate keyword density for web pages ranges from 3% to 8%. Consequently, Shahzad(Shahzad et al., 2020) et al. proposed a methodology for identifying spam, which includes a step to check whether the keyword density of a webpage exceeds a certain threshold; if it does, the webpage is excluded as spam. Many search engines and SEO tools classify webpages with keyword density exceeding a certain threshold as spam or compare the keyword density of the target webpage with that of other webpages to determined whether the target webpage is spam. Semrush  [^4] https://www.semrush.com/ compares the webpage’s keyword densities against top-ranking rivals’ keyword densities to see whether there is a disparity.
Many keyword stuffing practices concentrate keywords in specific sections of webpage content (e.g., title tags, anchor text, etc.). Therefore, SEO detection necessitates examining the distribution of keyword density within the content. Tan (Tan et al., 2018) et al. utilized a sliding window approach to derive local word density features in emails, which served as the basis for spam detection. We employed a sliding window technique to analyze keyword density across different sections of the text. If an abnormally high keyword density is detected within any window, it suggests that the document likely contains keyword stuffing.
Given that the RAG attack scenario involves manipulation during the search phase, similar to SEO, we also employed keyword density to detect the RAG attack. Following the method outlined in (Shahzad et al., 2020), we treated the terms in the user query as keywords and calculated the overall keyword density of the document. We calculated the overall keyword density and the maximum keyword density across all sliding windows for documents subjected to the baselines and FlippedRAG attacks, as well as for clean data. The sliding window sizes were set to 20, 50, and 100, with a step size of 5. The experimental results are presented in Table 9.
B.5.3. Mitigation Based on Paraphrasing
Paraphrasing is a widely utilized preprocessing defense originating from computer vision, where a generative model is employed to perform encoding and decoding operations on an adversarial image, thereby mitigating its harmful effects (Jain et al., 2023). In NLP, the paraphrasing defense can alter the text content, thereby changing the adversarial sequence or the attack target, which reduces the success rate of the attack on text models of language models.
Paraphrasing can be applied to both queries and documents. Paraphrasing applied to documents would require processing a large volume of documents, which is impractical in real-world scenarios. Therefore, we implemented paraphrasing on queries. Given a user query $q$, the paraphrasing defense leverages the LLM to rewrite $q$ into $q^{\prime}$, preserving the original semantics while ensuring that $q^{\prime}$ differs as much as possible from $q$. This approach aims to hinder the RAG attack by disrupting the retrieval ranking of poisoned documents associated with specific $q$. We use the following prompt to paraphrase the user query: “Paraphrase the following query and do not to use the words in the original query.”
B.5.4. Mitigation Based on Randomized Mask Smoothing
Randomized smoothing is a defense method aiming at achieving certified robustness against adversarial examples. It constructs a smoothed composite function by introducing random variables to the original function (model), thereby enabling the certification of its robustness under certain conditions. In the field of computer vision(CV), Levine et al. (Levine and Feizi, 2020) proposed a certifiable defense based on randomized smoothing against patch attacks in image classification tasks. This approach involves generating numerous noisy versions of the input and aggregating their classification results to mitigate the impact of adversarial patches, thereby achieving accurate classification. For ranking models, Wu et al. (Wu et al., 2022a) defined the concept of Certified Top-K Robustness for text ranking tasks and developed a smoothed ranking model based on random word substitution to address word substitution ranking attacks, thereby establishing a certifiably robust defense CertDR. CertDR relies on knowledge of the lexicon used by attackers to perform word substitution attacks, whereas Xiang et al. (Xiang et al., 2021) and Zeng et al. (Zeng et al., 2023) explore the use of masking to achieve random smoothing in both CV and NLP domains.
Given that one of the objectives of RAG attacks is to target retrieval ranking models, we explored the use of randomized mask smoothing as a defense mechanism to counteract RAG attacks.
B.5.5. Mitigation Based on RobustRAG
RobustRAG designs keyword-based and decoding-based algorithms for aggregating LLM responses. The framework first processes each retrieved passages constituting the context by combining it with the user query and inputs them individually into the LLM to generate multiple isolated responses. The keyword-based algorithm extracts keywords from these isolated responses and compiles them into a keyword set. Subsequently, the LLM generates the final response to the query based on the extracted keywords with large counts. The decoding-based algorithm aggregates the next-token probability vectors predicted from different isolated passages within the context and generates each subsequent token based on the aggregated probability vector, iteratively producing the final response. Experiments demonstrate that this voting-inspired defense mechanism proves particularly effective against poisoning attacks on factoid questions, consistently limiting the attack success rate to approximately 5% or lower.
For the mitigation approach discussed in this section, we exclusively adopt the keyword-based algorithm, as the decoding-based algorithm exhibits prohibitively long execution times, rendering it currently impractical for real-world deployment.